### v1

# Periodic Asynchrony: An Effective Method for Accelerating On-Policy Reinforcement Learning

Abstract
Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention, with growing efforts to reproduce and apply it. However, training efficiency remains a critical challenge. In mainstream RL frameworks, inference and training are typically deployed on the same devices. While this approach reduces costs through resource consolidation, its synchronous execution imposes a computational coupling that prevents concurrent inference and training. In this study, we are returning to the strategy of separating inference and training deployment, and by introducing improvements in the data loader, we transform the conventional synchronous architecture into a periodically asynchronous framework, which allows for demand-driven, independent, and elastic scaling of each component, while the accuracy of the algorithm remains completely equivalent to the synchronization method, with both belonging to the on-policy strategy. It is worth emphasizing that we apply a unified tri-model architecture in the training phase, and we also proposed a shared-prompt attention mask to reduce repetitive computation. In practice, these works have achieved at least a threefold overall performance improvement in RL training on NPU platforms, indicating its potential for widespread application.
1 Introduction
In recent years, with the rapid development of large language models (LLMs), Reinforcement learning (RL) has once again become a key technique for post-training and human alignment. In particular, DeepSeek-R1 [guo2025deepseek] introduced the Group Relative Policy Optimization (GRPO) algorithm, which successfully incentivized LLMs to exhibit stronger reasoning abilities in complex tasks. This work highlighted the potential of RL in improving reasoning and alignment in large models, and it has further stimulated widespread interest in efficient RL training methods across both academia and industry.
Nevertheless, despite the remarkable potential of RL in LLM alignment and reasoning tasks, training efficiency remains a critical challenge. This stems from the inherent complexity of the RL training pipeline. First, even the forward computation alone involves the primary training model (policy), an old policy model, and a reference model, while certain samples may additionally require a value model and a reward model [schulman2017proximal], all of which substantially increase computational overhead. Second, before each training step, the latest weights of the policy model are often used to generate a large number of chain-of-thought (CoT) [wei2022chain] trajectories to support optimization, further amplifying the inference burden. These factors not only constrain training throughput but also result in significant memory consumption, making efficiency improvements particularly challenging in large-scale model settings.
The optimization of reinforcement learning training has evolved continuously over time. Early DeepSpeed-Chat [yao2023dschat] primarily relied on ZeRO for distributed training, where the policy model itself was directly responsible for inference. This approach suffered from relatively low inference efficiency, which limited overall training throughput. Subsequently, an architecture that decouples inference from training is introduced, utilizing specialized inference engines such as vLLM [kwon2023efficient] to generate responses. Although these approaches increased resource usage slightly, they significantly improved inference efficiency, thereby substantially enhancing overall training performance, as exemplified by Open-R1 [openr1]. To further leverage computing resources and reduce idle time between training and inference, the industry introduced training–inference co-location schemes. These approaches are conceptually still based on training–inference separation but deploy both tasks within the same resource pool, sharing resources through repeated task switching. Typical instances include OpenRLHF [hu2024openrlhf]. However, such architecture remain tightly coupled, limiting independent scalability and preventing training and inference from executing simultaneously. Beyond optimizing inference, additional exploration has focused on the training frameworks themselves. For example, MindSpeed-RL [feng2025mindspeedrldistributeddataflow] and VERL [sheng2024hybridflow] have adopted Megatron-style 3D parallelism [megatron-lm] to train the policy model, improving throughput and enabling the training of larger models.
Against this backdrop, we revisit the inference–training separation strategy and enhance a conventional synchronous framework with new optimizations. Our approach builds on a Megatron-style 3D parallel architecture, further refined for improved training efficiency. Specifically, we add a background process in the data-loading module that acts as a producer, responsible for sending a full batch of prompts to the inferrence service. The main process acts as a consumer, immediately retrieving the outputs for training, where the received samples belonging to the same group are further divided into multiple micro-batches and trained sequentially. Once the entire batch has been consumed, the policy model’s weights are synchronized to the rollout workers, after which the next batch of asynchronous training begins. We refer to this mechanism as periodic asynchrony, which remains a purely on-policy asynchronous strategy and can thus be applied to any on-policy RL algorithm without modification. Benefiting from the decoupling of inference and training, the number of inference instances can be flexibly scaled according to inference efficiency, preventing inference from becoming the bottleneck. Furthermore, in GRPO, where each group of samples shares the same prompt, we propose a shared-prompt attention mask to reduce redundant computation and further improve efficiency. Experimental results demonstrate that, using 16 NPUs to train an 8B dense model with a 16k context length, our method achieves more than a threefold improvement in end-to-end RL training throughput.
2 Related Work
Improving the efficiency of reinforcement learning (RL) for large language models (LLMs) has been the subject of growing attention in both academia and industry. While synchronous frameworks dominate existing large-scale RL systems due to their algorithmic stability, they often suffer from underutilization of computational resources. To overcome these limitations, recent studies have explored asynchronous training paradigms, which aim to decouple inference from training and thereby enhance throughput and scalability. In this section, we review representative efforts in this direction, with a particular emphasis on asynchronous reinforcement learning approaches.
One of the earliest attempts in this direction is Asynchronous RLHF [noukhovitch2024asynchronous], which decouples the generation and learning processes, allowing new sample generation and old sample training to proceed in parallel. Its core idea is online but off-policy RLHF, where the model is updated using samples from previous iterations. While effective, this represents a typical off-policy strategy that may not directly apply to on-policy algorithms such as GRPO, and it remains constrained to relatively short-context settings.
Most recently, AReaL [fu2025areallargescaleasynchronousreinforcement] advanced the application of asynchronous RL to LLM reasoning tasks. After we had already developed the methods presented in this paper, we became aware of AReaL, which fully decouples generation from training while introducing a parameter $\eta$ to control data staleness. To better handle delayed samples, it incorporates a staleness-enhanced PPO variant, effectively modifying the on-policy paradigm to tolerate partially off-policy data. Additionally, its training start time depends on the completion of the first inference batch. Similarly, during the preparation of this manuscript, ROLL Flash [lu2025iirollflash] introduced a study sharing conceptual similarities with AReaL’s work. It also decouples training from inference and controls the proportion of off-policy data through an async ratio. Both AReaL and ROLL Flash are off-policy asynchronous frameworks that improve training efficiency but require modifications to standard on-policy RL algorithms to handle data bias. This design trades a small portion of data originally intended for strictly on-policy training to avoid inference-service blocking, effectively shifting it toward off-policy learning. Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
However, they all fall under off-policy asynchronous frameworks, requiring control of data bias and modifications to algorithms to maintain stability for on-policy model training. Distinct from these approaches, our method introduces a new perspective with several key innovations:
  1. 1.
    Periodic asynchrony, which guarantees purely on-policy asynchronous acceleration without compromising algorithmic stability.
  2. 2.
    Decoupling of prompt-level inference production and sample-level training consumption, which maximally overlaps inference and training.
  3. 3.
    A group-mask design that enables effective reuse of forward and backward computations for prompt tokens, thereby reducing redundant computation.
  4. 4.
    A distribution-consistent design across the policy model, old-policy model, and reference model, allowing more flexible scaling and simplifying the overall system.
  5. 5.
    Independent scalability of inference and training instances, ensuring that neither component becomes a performance bottleneck.
Together, these designs enable efficient long-context training with substantially fewer computational resources. All the components have been implemented in the project release v5.0.5 on September 5, 2025 (https://github.com/janelu9/EasyLLM).
3 Background
Training RL with Micro-Batching
It is well known that when using distributed architectures for large-scale model pre-training or fine-tuning, one typically does not train on an entire batch and immediately update the model parameters with the resulting gradients, as is common in single-GPU training for small models. A full batch usually contains hundreds or even thousands of samples, and directly processing it in large-model training can easily lead to GPU out-of-memory (OOM) errors. Instead, the batch is divided into multiple micro-batches, each containing only one or two samples. These micro-batches are processed sequentially, with a buffer tensor used to accumulate the gradients from each micro-step. The accumulated gradients are applied to update the model parameters only after all micro-batches within the batch have been processed. In practice, these two training strategies are mathematically equivalent.
We can formally demonstrate the equivalence of full-batch and micro-batch training by reviewing the gradient update rules. When training on an entire batch directly, the model parameters are updated as follows:
$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n}$ (1)
where $N$ is the total number of samples in the batch, and $g_{n}$ denotes the gradient computed from the $n$-th sample.
For micro-batch training, the update proceeds differently. First, the accumulated gradient is initialized. Then, for each micro-batch $j=1$ to $M$, the gradient is updated as
$g_{\text{acc}}:=g_{\text{acc}}+\frac{1}{m}\sum_{i=1}^{m}g_{i,j},$ (2)
where $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch. Finally, after processing all micro-batches, the model parameters are updated using the accumulated gradient:
$W_{t+1}=W_{t}-\alpha\frac{1}{M}g_{\text{acc}}.$ (3)
This formulation demonstrates that, mathematically, full-batch training and micro-batch training are equivalent. However, the micro-batch approach requires significantly less GPU memory. In particular, this delayed parameter update scheme is more suitable for reinforcement learning.
Consider the objective function of GRPO for generating responses from a single prompt:
$J_{GRPO}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{old},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{ref})\Big),$ (4)
where
$L_{i}(\pi_{\theta}\|\pi_{old},A_{i})=\min\Bigg(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},\ \text{clip}\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\Big)A_{i}\Bigg),$ (5)
$D_{KL}^{i}(\pi_{\theta}\|\pi_{ref})=\frac{\pi_{ref}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\frac{\pi_{ref}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (6)
For a batch of $N$ prompts, the objective function becomes:
$J_{batch}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{GRPO}^{i}$ (7)
$=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{G}\sum_{j=1}^{G}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big)$
$=\frac{1}{NG}\sum_{i=1}^{NG}\Big(L_{i}-\beta D_{KL}^{i}\Big).$
Let
$M=\frac{NG}{m}.$ (8)
Then, the batch can be trained using micro-batches as:
$J_{batch}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (9)
This demonstrates that reinforcement learning can also be trained using micro-batches. Compared with pre-training or fine-tuning, a batch now contains $NG$ samples, but GPU memory usage is greatly reduced. Consequently, reinforcement learning can handle sequences as long as those supported during fine-tuning.
4 System Design
4.1 Unified Tri-Model Architecture
When performing reinforcement learning with large-scale models using different frameworks for inference and training—for instance, vLLM for inference and Megatron for training—the process generally involves three main steps:
  1. 1.
    Synchronize the weights of the policy model to the inference engine.
  2. 2.
    The inference engine retrieves prompts from the dataloader and generates responses. A reward module evaluates the generated responses based on the prompts, distinguishing high-quality responses from lower-quality ones.
  3. 3.
    All generated samples and their associated scores are sent to the training engine for loss computation and backpropagation, updating the weights of the policy model.
For the third step, mainstream reinforcement learning systems typically require the computation of three types of logits in the forward pass: policy, old policy, and reference. These models share the same architecture, differing only in their weights. In this work, the training phase is implemented using a Megatron-like architecture, where a complete policy model is distributed across multiple computational units according to tensor parallelism and pipeline parallelism scales. During the forward computation of the old policy and reference models, each sub-network managed by the policy model is replicated twice across all computational units—once for computing the old policy logits and once for computing the reference logits. The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model. The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure1.
Figure 1: Architecture of unified triple models with shared distribution.
4.2 Shared-Prompt Attention
In reinforcement learning training, each group of samples shares the same prompt. In fact, during micro-batch training, if the micro-batch size is greater than 1, the prompt can be reused. When the prompt contains many tokens while the generated responses are relatively short, reusing the prompt can significantly reduce redundant computation and GPU memory usage, thereby improving overall training efficiency.
Compared with standard training, when adopting the shared-prompt approach, the following adjustments are required:
  1. 1.
    Concatenate the shared prompt with the token IDs of multiple responses. Correspondingly, the labels used for computing probabilities should also be concatenated, except that the prompt portion does not need to be included in the labels.
    $\left\{\begin{array}[]{l}x=\left[{{x_{p}},{x_{{r_{1}}}},{x_{{r_{2}}}}}\right]\\ y=\left[{{y_{{r_{1}}}},{y_{{r_{2}}}}}\right]\end{array}\right$. (10)
  2. 2.
    The position IDs need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
    $p=\left({0,1,...,\left|{{x_{p}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{1}}}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{2}}}}}\right|-1}\right)$ (11)
  3. 3.
    In the attention module, use a group mask as illustrated in the Figure2, instead of the standard causal mask. This ensures that the value vectors of the second and subsequent responses are only aggregated with the prompt’s value vectors, and not with the value vectors of the preceding responses.
    Figure 2: Shared-prompt attention mask.
    $\left\{\begin{array}[]{l}\hat{y}={\mathop{\rm model}\nolimits}\left({x,mas{k_{group}}}\right)\\ {{\hat{y}}^{old}}={{\mathop{\rm model}\nolimits}^{old}}\left({x,mas{k_{group}}}\right)\\ {{\hat{y}}^{ref}}={{\mathop{\rm model}\nolimits}^{ref}}\left({x,mas{k_{group}}}\right)\end{array}\right$. (12)
  4. 4.
    When computing probabilities, truncate the prompt tokens and retain only the response portion. Finally, the probabilities output by the three types of models, together with the reward scores, are fed into the GRPO objective function to compute the loss and perform iterative training.
    $\left\{\begin{array}[]{l}\pi=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}_{\left|{{x_{p}}}\right|:\left|x\right|}},y}\right)\\ {\pi_{old}}=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}^{old}}_{\left|{{x_{p}}}\right|:\left|x\right|},y}\right)\\ {\pi_{ref}}=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}^{ref}}_{\left|{{x_{p}}}\right|:\left|x\right|},y}\right)\\ Loss={J_{GRPO}}\left({\pi,{\pi_{old}},{\pi_{ref}},A}\right)\end{array}\right$. (13)
All other aspects remain the same as in standard training. This grouped micro-batch approach is generally more computationally efficient than simply concatenating micro-batches along the first axis, because the prompt tokens only need to be computed once, and padding is not required for shorter samples within a micro-batch. The advantage becomes especially pronounced when the prompt is long and the responses are relatively short.
4.3 Asynchronous On-Policy Reinforcement Learning
Asynchronous execution of inference and training is a key factor in accelerating reinforcement learning systems with a separated training-inference architecture. In such asynchronous systems, the critical aspect is minimizing the time the training engine waits for the rollout workers to return generated samples, i.e., the startup latency. In this work, we propose a simple approach that introduces a temporary data generator between the data loader and the trainer. This design transforms on-policy reinforcement learning from synchronous to asynchronous execution, significantly improving training efficiency without requiring any modifications to the underlying reinforcement learning algorithm. The reinforcement learning workflow incorporating the temporary sample generator is illustrated in Figure 3. It follows a typical producer–consumer pattern and consists of four main components:
  1. 1.
    Data source, which is the original data loader.
  2. 2.
    Temporary data generator, the additional component introduced in this work, which includes a background thread acting as a producer and a queue shared with the main process.
  3. 3.
    Inference service, responsible for receiving prompts and generating responses.
  4. 4.
    Trainer, which consumes the samples from the queue for training.
Figure 3: How our asynchronous system operates.
The core of our asynchronous design lies in the operation of the temporary generator. Within the temporary generator, a batch of prompts required for one training step is first retrieved from the standard data loader. A background thread is then launched to act as a producer. Inside this thread, a coroutine is used to concurrently and evenly dispatch the entire batch of prompts to multiple inference instances within the inference service. Upon receiving a large number of requests, the inference instances efficiently generate responses through continuous batching and return the completed responses back to the coroutine.
To fully exploit coroutine-level parallelism, the reward function is integrated directly into the coroutine. Each response returned by an inference instance is immediately scored, and once both inference and evaluation for a coroutine’s assignment are finished, the resulting responses and reward scores are placed into a queue shared with the main process. This queue serves as a buffer that temporarily stores the generated responses from inference service.
The main process, acting as the consumer, continuously retrieves completed samples from the queue and feeds them to the three models for training. Under this design, the start time of training is determined by the earliest completed sample among all prompts in the batch. Once all prompts in the batch have been processed, the inference service enters an idle state until the training step finishes and the updated policy weights are synchronized back to the rollout workers. Only then does the next training cycle begin. Therefore, we refer to this design as periodic asynchrony.
In this way, the system constitutes a fully on-policy asynchronous reinforcement learning framework that remains decoupled from the underlying RL algorithm, meaning that no algorithmic modifications are required to ensure correct operation within this asynchronous setup. Algorithm 1 is the pseudocode for the asynchronous reinforcement learning system proposed in this paper.
Algorithm 1 Periodic Asynchronous On-Policy Reinforcement Learning Algorithm
1:Dataloader $D$, training iterations $T$, batch size $B$
2:Trained policy model
3:Initialize shared queue $Q$ and gradient accumulator $O$
4:for $t=1$ to $T$ do
5:  Synchronize policy model weights to the rollout workers
6:  Retrieve prompts $P=\{p_{1},\dots,p_{B}\}$ from $D$
7:  Launch thread Producer:
8:   Create $B$ concurrent coroutine tasks
9:   Run in parallel for each $i=1,\dots,B$:
10:    $r_{i}\leftarrow$ InferenceService($p_{i}$)
11:    $a_{i}\leftarrow$ RewardFunction($r_{i}$)
12:    Enqueue $(a_{i},r_{i})$ into $Q$
13:  $O\leftarrow 0$
14:  $consumed\_count\leftarrow 0$
15:  while consumed_count<Bconsumed\_count<B do
16:     $(a_{i},r_{i})\leftarrow$ Dequeue($Q$)
17:     $s_{i}\leftarrow$ ProcessSample($p_{i},r_{i}$)
18:     Compute loss $L(s_{i})$ then backward
19:     $O\leftarrow O+\nabla L(s_{i})$
20:     $consumed\_count\leftarrow consumed\_count+1$
21:  end while
22:  Move policy’s weights into old-policy’s state_dict
23:  Update policy’s weights using gradient $O$
24:end for
(a) Training one step in a synchronous system.
(b) Training two steps in our asynchronous system.
Figure 4: Comparison of training steps in synchronous and asynchronous systems.
Figure 4(a) and Figure 4(b) illustrate the differences between the asynchronous training system studied in this paper and a synchronous training system. It is evident that in a synchronous system, the training process must wait for all prompts to be processed by the inference service. If the batch size is large, multiple inference batches may be required to complete all the prompts. The startup time for training is determined by the inference batches and the sum of the inference times of the slowest samples in each batch. Assuming the inference engine has a batch size of $m$ and the prompts of one training batch are divided into $n$ inference batches, the training startup time can be expressed as:
${T_{sync}}=\sum\limits_{j=0}^{n}{\max\left\{{t_{i}^{j}|i\in\left\{{0,1,\ldots,m}\right\}}\right\}}$ (14)
Where $t_{j}^{i}$ denotes the time required for the $i$-th prompt in the $j$-th batch to complete inference. In contrast, for an asynchronous system, the training startup time depends only on the earliest-completed prompt in the first inference batch, which can be expressed as:
$T_{\text{async}}=\min\left\{t_{i}^{0}\mid i\in\{0,1,\ldots,m\}\right\}$ (15)
Then, all samples in this batch are trained sequentially based on the order in which they enter the queue. Therefore, theoretically, if the training prompts in a batch cannot be processed by a single inference batch, the asynchronous system can achieve more than twice the speed of the synchronous system.
5 Implementation
The implementation in this study is based on the PyTorch training framework, with the training component using a 3D-parallel distributed architecture. Parts of the code from Megatron-Core v0.8.0 are incorporated.On the NPU platform, additional components from MindSpeed v2.1.0_core_r0.8.0, torch_npu, and several optimized operators included in torch_npu—such as npu_fusion_attention, an accelerated attention kernel supporting custom masks, which is designed as a counterpart to flash attention [dao2023flashattention]—are also employed.
For the inference component, vLLM is used as the engine, and on the NPU platform, ascend_vLLM 0.9.2.rc1 is adopted.
As in fine-tuning, all samples in the reinforcement learning training process are computed with their native sequence lengths for both forward and backward passes, without any additional padding; that is, the system uses dynamic-length training.
6 Experiments
This study primarily conducts comparative experiments on mathematical reasoning tasks. The reinforcement learning frameworks included in the comparison are: MindSpeed-RL, which adopts a training–inference shared-resource architecture; the synchronous reinforcement learning version implemented in this work under a training–inference decoupled design; and the asynchronous reinforcement learning version proposed in this study.
The main objective is to experimentally validate the performance between periodic asynchronous and synchronous training, as well as to compare the proposed periodic asynchronous system with other mainstream synchronous reinforcement learning frameworks on the same platform, while also evaluating the scalability of our periodic asynchronous system.
The primary evaluation metric is the end-to-end training throughput, such as the average number of tokens trained per second per device (TPSPD). This is calculated as the total number of tokens processed over several steps divided by the total elapsed time and then divided by the number of devices used.
Since no modifications were made to the reinforcement learning algorithm, accuracy metrics are reported only for reference, primarily to verify the correctness of algorithm reproduction.
6.1 Model and Environment Configuration
The models compared in this section are Qwen3-8B  [qwen3technicalreport] and DeepSeek-R1-Distill-Qwen-32B  [deepseekai2025deepseekr1incentivizingreasoningcapability] (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.). The training data comes from the publicly available math problem dataset DeepScaler [deepscaler2025]. All reinforcement learning experiments use the GRPO algorithm. The accuracy test set is the AIME24 [aime24].
6.2 Comparison between Asynchronous and Synchronous Training
In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k. A total of 16 NPUs were used. For the synchronous and asynchronous training configurations implemented in this work under the training–inference decoupled architecture, 8 NPUs were allocated to launch a single training instance with tensor parallelism ($\text{TP}=8$) and pipeline parallelism ($\text{PP}=1$). The remaining 8 NPUs were used to launch four inference instances with $\text{TP}=2$. For MindSpeed-RL, the training configuration used 16 NPUs to launch two training instances ($\text{TP}=8$, $\text{PP}=1$), while all 16 NPUs were also used to launch eight inference instances ($\text{TP}=2$).
Table 1: Accuracy and training efficiency comparison results on Qwen3-8B. Each framework was trained for 100 steps, and the total time consumed and the total number of tokens trained were extracted from the logs.
Framework GRPO Steps AIME24 Trained Tokens Training Hours NPUs Tokens/sec/device
Baseline 0 0.725 0 0 0 -
MindSpeed-RL 100 0.733 706.355M 198.942 16 61.641
Sync (our) 100 0.746 827.306M 143.678 16 99.966
Async (our) 100 0.758 917.609M 82.861 16 192.259
In terms of TPSPD, the training speed ratio among the three implementations was 1:1.62:3.12. For accuracy evaluation, we adopted a rule-based method, where the final answer predicted by the policy model is considered correct if it can be correctly extracted and matches the ground-truth answer in the dataset; otherwise, it is considered incorrect. The tests were conducted using the inference engine vLLM 0.9.1.rc, with an output length limit of 32k. Each test sample was evaluated 8 times, with inference parameters set to temperature = 0.7, top_p = 0.95, and top_k = 20. In total, 240 test problems were evaluated. It is evident that the models trained by the three frameworks achieved similar accuracy on the AIME24 benchmark, with variations within normal random fluctuations, supporting the equivalence in effectiveness between the asynchronous and synchronous methods implemented in this work.
Figure 5: Training step-wise average reward score comparison
We also monitored the reward scores throughout the training process of the three frameworks, as depicted in Figure 5. The reward score trajectories of our synchronous and asynchronous methods are nearly indistinguishable, further reinforcing the assertion that the proposed asynchronous approach is computationally equivalent to its synchronous counterpart. Both curves show a slight upward trend, which can likely be attributed to the baseline model having already been pre-trained on similar data. Notably, in the initial phases of training, the reward mean curve for Mindspeed-RL diverges significantly from that of our approach. This disparity may arise from differences in system architecture. Additionally, it is important to note that, in contrast to our environment, Mindspeed-RL utilizes vLLM version 0.9.1 and vllm_ascend version 0.9.0rc2.
The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments are conducted. For our asynchronous framework, the configuration is as follows: 32 NPUs are used to launch a single training instance with $\text{TP}=8$ and $\text{PP}=4$, while the remaining 16 NPUs are used to launch four inference instances with $\text{TP}=4$. In contrast, for the Mindspeed-RL configuration, 64 NPUs are used to launch two training instances with $\text{TP}=8$ and $\text{PP}=4$, and simultaneously, the same 64 NPUs are used to launch 16 inference instances with $\text{TP}=4$.
Table 2: Accuracy and training efficiency comparison results on DeepSeek-R1-Distill-Qwen-32B. Each framework was trained for 20 steps, and the total time consumed and the total number of tokens trained were extracted from the logs.
Framework GRPO Steps AIME24 Trained Tokens Training Hours NPUs Tokens/sec/device
Baseline 0 0.726 0 0 0 -
MindSpeed-RL 20 0.679 105.678M 69.116 64 6.636
Async (our) 20 0.737 123.999M 21.453 48 33.449
The evaluation methodology is identical to that used in the first set of experiments. According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is 1 : 5.05.
Combining the results of the two experimental groups, we obtain a clear comparative plot of the training speeds, as shown in Figure 6.
Figure 6: TPSPD between MindSpeed-RL and Async(ours)
6.3 Asynchronous Scalability Experiments
We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. Under an inter-node bandwidth of 100 Gb/s, the results are shown in Table 3.
Table 3: Scalability comparison results. TPSPD performance of training Qwen3-8B on 16, 32, and 64 NPUs using our asynchronous framework.
Model NPUs Training Instance Inference Instance Training Steps Tokens/sec/device
Qwen3-8B 16 1 4 20 188.162
32 2 8 20 171.824
64 4 16 20 163.208
When expressed in TPS (tokens per second), the results are more intuitively illustrated in Figure 7. Training with 32 NPUs shows a 1.83× speedup compared to 16 NPUs, while using 64 NPUs results in a 1.9× speedup over 32 NPUs. Overall, the proposed asynchronous reinforcement learning framework demonstrates near-linear scaling in training speed as the number of devices increases.
Figure 7: TPS of our framework under different scales of data parallelism
7 Conclusion
This paper presents an asynchronous on-policy reinforcement learning implementation that maintains the same training stability as synchronous RL methods. The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated. Prompt-level asynchronous parallelism maximizes the overlap between inference and training. During training, a three-mode (policy, old policy, reference) unified and identically distributed architecture is introduced, along with a group-mask attention mechanism that provides clear advantages when prompts are long and responses are short.
Experiments show that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform. The results also support the equivalence in training effectiveness between the asynchronous and synchronous methods. Furthermore, the decoupling of training and inference allows both to be scaled independently and flexibly, achieving optimal balance and maximizing the speed-to-device ratio. The decoupling of training architecture and algorithm also implies that the proposed framework can be applied to other on-policy reinforcement learning algorithms without any modifications for compatibility with asynchronous training architectures.

### v2

@@
-Periodic Asynchrony: An Effective Method for Accelerating On-Policy Reinforcement Learning
+Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning
 Abstract
@@
 The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model.
+Figure 1: Architecture of unified triple models with shared distribution.
 The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure 1.
-Figure 1: Architecture of unified triple models with shared distribution.
+The three types of models share an identical architecture and are distributed in the same parallel layout.
+This design allows their logits to be computed simultaneously without requiring any additional topology for the reference model.
+The inference stage, in contrast, is handled by an independent inference engine such as vLLM.
+Before training each batch, the weights of the policy model are synchronized to all inference engines.
+The inference engines then receive prompts from the dataset, generate responses, and return them for scoring.
+Each scored response is concatenated with its corresponding prompt to form a training sample, which is fed into all three models for forward computation.
+In a single micro-step, we obtain the logits from the policy, old policy, and reference models simultaneously.
+After all samples in a batch complete their forward passes—and before the policy model’s weights are updated—its current weights are moved into the old policy model’s state_dict.
 4.2 Shared-Prompt Attention
@@
   2. 2.
-    The position IDs need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
+    The position indices need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
     $p=\left({0,1,...,\left|{{x_{p}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{1}}}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{2}}}}}\right|-1}\right)$ (11)
@@
 The advantage becomes especially pronounced when the prompt is long and the responses are relatively short.
-4.3 Asynchronous On-Policy Reinforcement Learning
+4.3 Periodic Asynchronous Reinforcement Learning
 Asynchronous execution of inference and training is a key factor in accelerating reinforcement learning systems with a separated training-inference architecture.
@@
 Algorithm 1 is the pseudocode for the asynchronous reinforcement learning system proposed in this paper.
-Algorithm 1 Periodic Asynchronous On-Policy Reinforcement Learning Algorithm
+Algorithm 1 Periodic Asynchronous Reinforcement Learning Algorithm
 1:Dataloader $D$, training iterations $T$, batch size $B$
@@
 (b) Training two steps in our asynchronous system.
-Figure 4: Comparison of training steps in synchronous and asynchronous systems.
+Figure 4: Comparison of training steps in synchronous and our asynchronous systems.
 Figure 4(a) and Figure 4(b) illustrate the differences between the asynchronous training system studied in this paper and a synchronous training system.
@@
 $T_{\text{async}}=\min\left\{t_{i}^{0}\mid i\in\{0,1,\ldots,m\}\right\}$ (15)
-Then, all samples in this batch are trained sequentially based on the order in which they enter the queue.
-Therefore, theoretically, if the training prompts in a batch cannot be processed by a single inference batch, the asynchronous system can achieve more than twice the speed of the synchronous system.
+After that, all samples in the batch are trained sequentially in the order in which they enter the queue.
+It is clear that when inference speed and training speed are well balanced, the sum of their waiting times is minimized, resulting in optimal end-to-end training throughput on a given hardware configuration.
+The decoupled inference–training architecture naturally enables such balancing by allowing independent configuration of inference and training instances.
+Therefore, in theory, when a training batch cannot be processed within a single inference batch, the asynchronous system can achieve more than twice the startup speed of the synchronous system.
 5 Implementation
@@
 Since no modifications were made to the reinforcement learning algorithm, accuracy metrics are reported only for reference, primarily to verify the correctness of algorithm reproduction.
+The experiments in this study were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight NPUs, each providing 64 GB of memory.
+The intra-node interconnect bandwidth is 196 GB/s, while the inter-node bandwidth is 100 Gb/s.
 6.1 Model and Environment Configuration
@@
 MindSpeed-RL 100 0.733 706.355M 198.942 16 61.641
-Sync (our) 100 0.746 827.306M 143.678 16 99.966
-Async (our) 100 0.758 917.609M 82.861 16 192.259
+Sync (ours) 100 0.746 827.306M 143.678 16 99.966
+Async (ours) 100 0.758 917.609M 82.861 16 192.259
 In terms of TPSPD, the training speed ratio among the three implementations was 1:1.62:3.12.
@@
 MindSpeed-RL 20 0.679 105.678M 69.116 64 6.636
-Async (our) 20 0.737 123.999M 21.453 48 33.449
+Async (ours) 20 0.737 123.999M 21.453 48 33.449
 The evaluation methodology is identical to that used in the first set of experiments.
@@
 6.3 Asynchronous Scalability Experiments
-We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group.
-Under an inter-node bandwidth of 100 Gb/s, the results are shown in Table 3.
+We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. The results are shown in Table 3.
 Table 3: Scalability comparison results. TPSPD performance of training Qwen3-8B on 16, 32, and 64 NPUs using our asynchronous framework.
@@
 7 Conclusion
-This paper presents an asynchronous on-policy reinforcement learning implementation that maintains the same training stability as synchronous RL methods.
+This paper presents an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods.
 The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated.

### v3

@@
-Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning
+Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning for Large Language Models
 Abstract
@@
 1 Introduction
-In recent years, with the rapid development of large language models (LLMs), Reinforcement learning (RL) has once again become a key technique for post-training and human alignment.
-In particular, DeepSeek-R1 [guo2025deepseek] introduced the Group Relative Policy Optimization (GRPO) algorithm, which successfully incentivized LLMs to exhibit stronger reasoning abilities in complex tasks.
-This work highlighted the potential of RL in improving reasoning and alignment in large models, and it has further stimulated widespread interest in efficient RL training methods across both academia and industry.
-Nevertheless, despite the remarkable potential of RL in LLM alignment and reasoning tasks, training efficiency remains a critical challenge.
-This stems from the inherent complexity of the RL training pipeline.
-First, even the forward computation alone involves the primary training model (policy), an old policy model, and a reference model, while certain samples may additionally require a value model and a reward model [schulman2017proximal], all of which substantially increase computational overhead.
-Second, before each training step, the latest weights of the policy model are often used to generate a large number of chain-of-thought (CoT) [wei2022chain] trajectories to support optimization, further amplifying the inference burden.
-These factors not only constrain training throughput but also result in significant memory consumption, making efficiency improvements particularly challenging in large-scale model settings.
-The optimization of reinforcement learning training has evolved continuously over time.
-Early DeepSpeed-Chat [yao2023dschat] primarily relied on ZeRO for distributed training, where the policy model itself was directly responsible for inference.
-This approach suffered from relatively low inference efficiency, which limited overall training throughput.
-Subsequently, an architecture that decouples inference from training is introduced, utilizing specialized inference engines such as vLLM [kwon2023efficient] to generate responses.
-Although these approaches increased resource usage slightly, they significantly improved inference efficiency, thereby substantially enhancing overall training performance, as exemplified by Open-R1 [openr1].
-To further leverage computing resources and reduce idle time between training and inference, the industry introduced training–inference co-location schemes.
-These approaches are conceptually still based on training–inference separation but deploy both tasks within the same resource pool, sharing resources through repeated task switching.
-Typical instances include OpenRLHF [hu2024openrlhf].
-However, such architecture remain tightly coupled, limiting independent scalability and preventing training and inference from executing simultaneously.
-Beyond optimizing inference, additional exploration has focused on the training frameworks themselves.
-For example, MindSpeed-RL [feng2025mindspeedrldistributeddataflow] and VERL [sheng2024hybridflow] have adopted Megatron-style 3D parallelism [megatron-lm] to train the policy model, improving throughput and enabling the training of larger models.
-Against this backdrop, we revisit the inference–training separation strategy and enhance a conventional synchronous framework with new optimizations.
-Our approach builds on a Megatron-style 3D parallel architecture, further refined for improved training efficiency.
-Specifically, we add a background process in the data-loading module that acts as a producer, responsible for sending a full batch of prompts to the inferrence service.
-The main process acts as a consumer, immediately retrieving the outputs for training, where the received samples belonging to the same group are further divided into multiple micro-batches and trained sequentially.
-Once the entire batch has been consumed, the policy model’s weights are synchronized to the rollout workers, after which the next batch of asynchronous training begins.
-We refer to this mechanism as periodic asynchrony, which remains a purely on-policy asynchronous strategy and can thus be applied to any on-policy RL algorithm without modification.
-Benefiting from the decoupling of inference and training, the number of inference instances can be flexibly scaled according to inference efficiency, preventing inference from becoming the bottleneck.
-Furthermore, in GRPO, where each group of samples shares the same prompt, we propose a shared-prompt attention mask to reduce redundant computation and further improve efficiency.
-Experimental results demonstrate that, using 16 NPUs to train an 8B dense model with a 16k context length, our method achieves more than a threefold improvement in end-to-end RL training throughput.
+Reinforcement learning (RL) has re-emerged as a key technique for post-training and aligning large language models (LLMs).
+Following the introduction of GRPO by DeepSeek-R1 Guo et al.
+(2025), RL has demonstrated strong potential in improving reasoning capabilities, sparking growing interest in efficient RL pipelines across academia and industry.
+Despite these advances, RL for LLMs still faces severe efficiency challenges.
+The training pipeline requires multiple models in the forward pass—the policy model, old-policy model, and reference model, and in some cases a value and reward model Schulman et al.
+(2017)—leading to substantial computational overhead.
+Moreover, each training step typically depends on generating large numbers of chain-of-thought (CoT) Wei et al.
+(2022) trajectories using the latest policy weights, further increasing inference cost and memory usage.
+To improve efficiency, early systems such as DeepSpeed-Chat Yao et al.
+(2023) relied on ZeRO-based distributed training, where the policy model handled inference directly, resulting in limited throughput.
+Later approaches decoupled training and inference, leveraging inference engines like vLLM Kwon et al.
+(2023) to accelerate rollout generation, as demonstrated by Open-R1 Hugging Face (2025).
+Industry systems such as OpenRLHF Hu et al.
+(2024) push this further through training–inference co-location, although task switching still prevents full concurrency and independent scalability.
+Meanwhile, frameworks such as MindSpeed-RL Feng et al.
+(2025) and VERL Sheng et al.
+(2024) improve the training side through Megatron-style 3D parallelism Shoeybi et al.
+(2019).
+Motivated by these limitations, we revisit the training–inference separation strategy and introduce a synchronous framework enhanced with new system-level optimizations.
+Our design uses a Megatron-style 3D parallel backbone and adds a background producer process that continuously dispatches prompts to inference workers, while the main consumer process retrieves outputs and performs training.
+Samples from each group are split into micro-batches and trained sequentially, and model weights are synchronized only after the full batch is consumed.
+This periodic asynchrony achieves purely on-policy asynchronous acceleration and is compatible with any on-policy RL algorithm.
+The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation.
+Experiments show that on a 16-NPU setup training an 8B model with 16k context, our approach delivers over a threefold improvement in end-to-end RL throughput.
 2 Related Work
-Improving the efficiency of reinforcement learning (RL) for large language models (LLMs) has been the subject of growing attention in both academia and industry.
-While synchronous frameworks dominate existing large-scale RL systems due to their algorithmic stability, they often suffer from underutilization of computational resources.
-To overcome these limitations, recent studies have explored asynchronous training paradigms, which aim to decouple inference from training and thereby enhance throughput and scalability.
-In this section, we review representative efforts in this direction, with a particular emphasis on asynchronous reinforcement learning approaches.
-One of the earliest attempts in this direction is Asynchronous RLHF [noukhovitch2024asynchronous], which decouples the generation and learning processes, allowing new sample generation and old sample training to proceed in parallel.
-Its core idea is online but off-policy RLHF, where the model is updated using samples from previous iterations.
-While effective, this represents a typical off-policy strategy that may not directly apply to on-policy algorithms such as GRPO, and it remains constrained to relatively short-context settings.
-Most recently, AReaL [fu2025areallargescaleasynchronousreinforcement] advanced the application of asynchronous RL to LLM reasoning tasks.
-After we had already developed the methods presented in this paper, we became aware of AReaL, which fully decouples generation from training while introducing a parameter $\eta$ to control data staleness.
+Improving the efficiency of reinforcement learning (RL) for large language models (LLMs) has attracted increasing attention across academia and industry.
+Although synchronous frameworks dominate large-scale RL systems due to their stability, they often underutilize computational resources.
+To address this, recent studies explore asynchronous training paradigms that decouple inference from training to enhance throughput and scalability.
+We review representative efforts, with emphasis on asynchronous RL methods.
+An early attempt is Asynchronous RLHF Noukhovitch et al.
+(2024), which decouples generation and learning so that new sample generation and old-sample training proceed in parallel.
+Its key idea is online but off-policy RLHF, updating models with samples from previous iterations.
+While effective, it remains a typical off-policy strategy and is less suitable for strictly on-policy algorithms such as GRPO, especially in long-context settings.
+Most recently, AReaL Fu et al.
+(2025) advanced the application of asynchronous RL to LLM reasoning tasks by fully decoupling generation from training.
+It introduces a parameter $\eta$ to control data staleness, enabling better handling of delayed samples.
 To better handle delayed samples, it incorporates a staleness-enhanced PPO variant, effectively modifying the on-policy paradigm to tolerate partially off-policy data.
 Additionally, its training start time depends on the completion of the first inference batch.
-Similarly, during the preparation of this manuscript, ROLL Flash [lu2025iirollflash] introduced a study sharing conceptual similarities with AReaL’s work.
-It also decouples training from inference and controls the proportion of off-policy data through an async ratio.
+In a similar vein, ROLL Flash Lu et al.
+(2025), also recently introduced, shares conceptual similarities with AReaL.
+Like AReaL, it decouples training from inference and controls the proportion of off-policy data through an async ratio.
 Both AReaL and ROLL Flash are off-policy asynchronous frameworks that improve training efficiency but require modifications to standard on-policy RL algorithms to handle data bias.
@@
 Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
-However, they all fall under off-policy asynchronous frameworks, requiring control of data bias and modifications to algorithms to maintain stability for on-policy model training.
-Distinct from these approaches, our method introduces a new perspective with several key innovations:
-  1. 1.
-    Periodic asynchrony, which guarantees purely on-policy asynchronous acceleration without compromising algorithmic stability.
-  2. 2.
-    Decoupling of prompt-level inference production and sample-level training consumption, which maximally overlaps inference and training.
-  3. 3.
-    A group-mask design that enables effective reuse of forward and backward computations for prompt tokens, thereby reducing redundant computation.
-  4. 4.
-    A distribution-consistent design across the policy model, old-policy model, and reference model, allowing more flexible scaling and simplifying the overall system.
-  5. 5.
-    Independent scalability of inference and training instances, ensuring that neither component becomes a performance bottleneck.
+In contrast, our method takes a different perspective with several innovations:
+  • Algorithm-decoupled periodic asynchrony enabling purely on-policy asynchronous acceleration without sacrificing stability.
+  • Decoupling prompt-level inference production from sample-level training consumption, maximizing overlap between the two.
+  • A shared-prompt mask for efficient reuse of forward/backward computation on prompt tokens.
+  • A distribution-consistent design across policy, old-policy, and reference models, simplifying scaling and system structure.
+  • Independent scalability of inference and training instances, preventing bottlenecks in either component.
 Together, these designs enable efficient long-context training with substantially fewer computational resources.
-All the components have been implemented in the project release v5.0.5 on September 5, 2025 (https://github.com/janelu9/EasyLLM).
+All components are implemented in project release v5.0.5 (September 5, 2025): https://github.com/janelu9/EasyLLM.
 3 Background
@@
 In practice, these two training strategies are mathematically equivalent.
-We can formally demonstrate the equivalence of full-batch and micro-batch training by reviewing the gradient update rules.
-When training on an entire batch directly, the model parameters are updated as follows:
-$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n}$ (1)
-where $N$ is the total number of samples in the batch, and $g_{n}$ denotes the gradient computed from the $n$-th sample.
-For micro-batch training, the update proceeds differently. First, the accumulated gradient is initialized. Then, for each micro-batch $j=1$ to $M$, the gradient is updated as
-$g_{\text{acc}}:=g_{\text{acc}}+\frac{1}{m}\sum_{i=1}^{m}g_{i,j},$ (2)
-where $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch.
-Finally, after processing all micro-batches, the model parameters are updated using the accumulated gradient:
-$W_{t+1}=W_{t}-\alpha\frac{1}{M}g_{\text{acc}}.$ (3)
-This formulation demonstrates that, mathematically, full-batch training and micro-batch training are equivalent.
-However, the micro-batch approach requires significantly less GPU memory.
-In particular, this delayed parameter update scheme is more suitable for reinforcement learning.
+The gradient update rule for full-batch training is:
+$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n},$ (1)
+and for micro-batch training, after accumulating gradients over all micro-batches, the update rule is:
+$W_{t+1}=W_{t}-\alpha\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}g_{i,j},$ (2)
+where $M$ is the total number of micro-batches, $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch.
+This demonstrates that both training strategies are mathematically equivalent, but micro-batching offers significant memory savings.
 Consider the objective function of GRPO for generating responses from a single prompt:
-$J_{GRPO}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{old},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{ref})\Big),$ (4)
-where
-$L_{i}(\pi_{\theta}\|\pi_{old},A_{i})=\min\Bigg(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},\ \text{clip}\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\varepsilon,1+\varepsilon\Big)A_{i}\Bigg),$ (5)
-$D_{KL}^{i}(\pi_{\theta}\|\pi_{ref})=\frac{\pi_{ref}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\frac{\pi_{ref}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (6)
-For a batch of $N$ prompts, the objective function becomes:
-$J_{batch}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{GRPO}^{i}$ (7)
+$J_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big).$ (3)
+The PPO-style clipped advantage term can be split into two lines as follows:
+$L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})=\min\Bigg($ $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)}A_{i},$
+$\text{clip}\!\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)},1-\varepsilon,\,1+\varepsilon\Big)A_{i}\Bigg).$ (4)
+The KL divergence term is given by:
+$D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})=\frac{\pi_{\mathrm{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\frac{\pi_{\mathrm{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (5)
+For a batch of $N$ prompts, the objective becomes:
+$J_{\mathrm{batch}}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{\mathrm{GRPO}}^{i}$
 $=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{G}\sum_{j=1}^{G}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big)$
-$=\frac{1}{NG}\sum_{i=1}^{NG}\Big(L_{i}-\beta D_{KL}^{i}\Big).$
-Let
+$=\frac{1}{NG}\sum_{i=1}^{NG}\Big(L_{i}-\beta D_{KL}^{i}\Big).$ (6)
+Define
 $M=\frac{NG}{m}.$ (7)
-Then, the batch can be trained using micro-batches as:
-$J_{batch}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (9)
+Then the batch objective can be rewritten using micro-batches:
+$J_{\mathrm{batch}}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (8)
 This demonstrates that reinforcement learning can also be trained using micro-batches.
@@
   1. 1.
-    Synchronize the weights of the policy model to the inference engine.
+    Synchronize the weights of the policy model with the inference engine.
   2. 2.
     The inference engine retrieves prompts from the dataloader and generates responses.
-A reward module evaluates the generated responses based on the prompts, distinguishing high-quality responses from lower-quality ones.
+A reward module evaluates these responses based on the corresponding prompts, distinguishing high-quality outputs from lower-quality ones.
   3. 3.
-    All generated samples and their associated scores are sent to the training engine for loss computation and backpropagation, updating the weights of the policy model.
+    The generated samples and their associated scores are then sent to the training engine, which computes the loss and performs backpropagation to update the policy model parameters.
 For the third step, mainstream reinforcement learning systems typically require the computation of three types of logits in the forward pass: policy, old policy, and reference.
@@
 The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model.
-Figure 1: Architecture of unified triple models with shared distribution.
+Figure 1: Triple Models with Shared Distribution.
 The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure 1.
@@
 When the prompt contains many tokens while the generated responses are relatively short, reusing the prompt can significantly reduce redundant computation and GPU memory usage, thereby improving overall training efficiency.
-Compared with standard training, when adopting the shared-prompt approach, the following adjustments are required:
+Compared with standard training, when adopting the shared-prompt approach, the following modifications are made:
   1. 1.
-    Concatenate the shared prompt with the token IDs of multiple responses.
-Correspondingly, the labels used for computing probabilities should also be concatenated, except that the prompt portion does not need to be included in the labels.
+    Concatenate the shared prompt with the token IDs of multiple responses. The corresponding labels are concatenated accordingly, while the prompt portion is excluded from the labels.
     $\left\{\begin{array}[]{l}x=\left[x_{p},x_{r_{1}},x_{r_{2}}\right]\\ y=\left[y_{r_{1}},y_{r_{2}}\right]\end{array}\right$. (9)
+    In this formulation, $x_{p}$ denotes the token IDs of the shared prompt, and $x_{r_{1}},x_{r_{2}}$ correspond to the tokens of the individual responses.
   2. 2.
-    The position indices need to be rearranged. For the second and subsequent responses, their starting positions should immediately follow the prompt, maintaining continuity with the prompt tokens.
-    $p=\left({0,1,...,\left|{{x_{p}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{1}}}}}\right|-1,\left|{{x_{p}}}\right|,...,\left|{{x_{p}}}\right|+\left|{{x_{{r_{2}}}}}\right|-1}\right)$ (11)
+    Rearrange the position indices such that each response starts immediately after the prompt, ensuring positional continuity with the shared prompt.
+    $p=($ $0,1,\dots,|x_{p}|-1,\ |x_{p}|,\dots,|x_{p}|+|x_{r_{1}}|-1,$
+    $|x_{p}|,\dots,|x_{p}|+|x_{r_{2}}|-1)$ (10)
+    Here, the notation $|\cdot|$ indicates sequence length.
   3. 3.
-    In the attention module, use a group mask as illustrated in the Figure 2, instead of the standard causal mask.
-This ensures that the value vectors of the second and subsequent responses are only aggregated with the prompt’s value vectors, and not with the value vectors of the preceding responses.
+    In the attention module, a shared-prompt mask (Figure 2) is used instead of the standard causal mask.
+This design restricts each response token to attend only to the shared prompt, preventing information leakage across responses.
     Figure 2: Shared-prompt attention mask.
-    $\left\{\begin{array}[]{l}\hat{y}={\mathop{\rm model}\nolimits}\left({x,mas{k_{group}}}\right)\\ {{\hat{y}}^{old}}={{\mathop{\rm model}\nolimits}^{old}}\left({x,mas{k_{group}}}\right)\\ {{\hat{y}}^{ref}}={{\mathop{\rm model}\nolimits}^{ref}}\left({x,mas{k_{group}}}\right)\end{array}\right$.
-(12)
   4. 4.
-    When computing probabilities, truncate the prompt tokens and retain only the response portion.
-Finally, the probabilities output by the three types of models, together with the reward scores, are fed into the GRPO objective function to compute the loss and perform iterative training.
-    $\left\{\begin{array}[]{l}\pi=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}_{\left|{{x_{p}}}\right|:\left|x\right|}},y}\right)\\ {\pi_{old}}=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}^{old}}_{\left|{{x_{p}}}\right|:\left|x\right|},y}\right)\\ {\pi_{ref}}=-{\mathop{\rm CrossEntropy}\nolimits}\left({{{\hat{y}}^{ref}}_{\left|{{x_{p}}}\right|:\left|x\right|},y}\right)\\ Loss={J_{GRPO}}\left({\pi,{\pi_{old}},{\pi_{ref}},A}\right)\end{array}\right$.
-(13)
+    For probability computation, the prompt tokens are discarded and the loss is computed only over the response tokens:
+    $\pi=-\mathop{\rm CrossEntropy}\nolimits\left(\hat{y}_{|x_{p}|:|x|},y\right),$ (11)
+    where $\hat{y}_{|x_{p}|:|x|}$ represents the predicted logits of the response tokens.
 All other aspects remain the same as in standard training.
@@
 This design transforms on-policy reinforcement learning from synchronous to asynchronous execution, significantly improving training efficiency without requiring any modifications to the underlying reinforcement learning algorithm.
-The reinforcement learning workflow incorporating the temporary sample generator is illustrated in Figure 3.
-It follows a typical producer–consumer pattern and consists of four main components:
+The reinforcement learning workflow incorporating the temporary sample generator is illustrated in Figure 3. It follows a typical producer–consumer pattern and consists of four main components:
   1. 1.
@@
     Trainer, which consumes the samples from the queue for training.
-Figure 3: How our asynchronous system operates.
+Figure 3: Operational Mechanism of the Asynchronous System.
 The core of our asynchronous design lies in the operation of the temporary generator.
@@
 Therefore, we refer to this design as periodic asynchrony.
+Algorithm 1 Periodic Asynchronous Reinforcement Learning Algorithm
+Input: Dataloader $D$, training iterations $T$, batch size $B$
+Output: Trained policy model
+1: Initialize shared queue $Q$ and gradient accumulator $O$
+2:for $t=1$ to $T$ do
+3:  Synchronize policy model weights to the rollout workers
+4:  Retrieve prompts $P=\{p_{1},\dots,p_{B}\}$ from $D$
+5:  Producer thread:
+6:     Create $B$ concurrent coroutine tasks
+7:     Run in parallel for each $i=1,\dots,B$:
+8:      $r_{i}\leftarrow$ InferenceService($p_{i}$)
+9:      $a_{i}\leftarrow$ RewardFunction($r_{i}$)
+10:      Enqueue $(a_{i},r_{i})$ into $Q$
+11:  $O\leftarrow 0$
+12:  $consumed\_count\leftarrow 0$
+13:  while consumed_count<Bconsumed\_count<B do
+14:     $(a_{i},r_{i})\leftarrow$ Dequeue($Q$)
+15:     $s_{i}\leftarrow$ ProcessSample($p_{i},r_{i}$)
+16:     Compute loss $L(s_{i})$ then backward
+17:     $O\leftarrow O+\nabla L(s_{i})$
+18:     $consumed\_count\leftarrow consumed\_count+1$
+19:  end while
+20:  Move policy’s weights into old-policy’s modules
+21:  Update policy’s weights using gradient $O$
+22:end for
 In this way, the system constitutes a fully on-policy asynchronous reinforcement learning framework that remains decoupled from the underlying RL algorithm, meaning that no algorithmic modifications are required to ensure correct operation within this asynchronous setup.
 Algorithm 1 is the pseudocode for the asynchronous reinforcement learning system proposed in this paper.
-Algorithm 1 Periodic Asynchronous Reinforcement Learning Algorithm
-1:Dataloader $D$, training iterations $T$, batch size $B$
-2:Trained policy model
-3:Initialize shared queue $Q$ and gradient accumulator $O$
-4:for $t=1$ to $T$ do
-5:  Synchronize policy model weights to the rollout workers
-6:  Retrieve prompts $P=\{p_{1},\dots,p_{B}\}$ from $D$
-7:  Launch thread Producer:
-8:   Create $B$ concurrent coroutine tasks
-9:   Run in parallel for each $i=1,\dots,B$:
-10:    $r_{i}\leftarrow$ InferenceService($p_{i}$)
-11:    $a_{i}\leftarrow$ RewardFunction($r_{i}$)
-12:    Enqueue $(a_{i},r_{i})$ into $Q$
-13:  $O\leftarrow 0$
-14:  $consumed\_count\leftarrow 0$
-15:  while consumed_count<Bconsumed\_count<B do
-16:     $(a_{i},r_{i})\leftarrow$ Dequeue($Q$)
-17:     $s_{i}\leftarrow$ ProcessSample($p_{i},r_{i}$)
-18:     Compute loss $L(s_{i})$ then backward
-19:     $O\leftarrow O+\nabla L(s_{i})$
-20:     $consumed\_count\leftarrow consumed\_count+1$
-21:  end while
-22:  Move policy’s weights into old-policy’s state_dict
-23:  Update policy’s weights using gradient $O$
-24:end for
-(a) Training one step in a synchronous system.
-(b) Training two steps in our asynchronous system.
-Figure 4: Comparison of training steps in synchronous and our asynchronous systems.
+(a) Synchronous training step.
+(b) Asynchronous two-step training.
+Figure 4: Comparison of synchronous vs. asynchronous training.
 Figure 4(a) and Figure 4(b) illustrate the differences between the asynchronous training system studied in this paper and a synchronous training system.
@@
 5 Implementation
-The implementation in this study is based on the PyTorch training framework, with the training component using a 3D-parallel distributed architecture.
-Parts of the code from Megatron-Core v0.8.0 are incorporated.On the NPU platform, additional components from MindSpeed v2.1.0_core_r0.8.0, torch_npu, and several optimized operators included in torch_npu—such as npu_fusion_attention, an accelerated attention kernel supporting custom masks, which is designed as a counterpart to flash attention [dao2023flashattention]—are also employed.
-For the inference component, vLLM is used as the engine, and on the NPU platform, ascend_vLLM 0.9.2.rc1 is adopted.
-As in fine-tuning, all samples in the reinforcement learning training process are computed with their native sequence lengths for both forward and backward passes, without any additional padding; that is, the system uses dynamic-length training.
+Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture.
+We incorporate parts of Megatron-Core v0.8.0 and DeepSpeed v0.17.2 Rasley et al.
+(2020).
+On the NPU platform, we additionally use MindSpeed v2.1.0_core_r0.8.0, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023).
+For inference, we use vLLM, with ascend_vLLM 0.9.2.rc1 on NPUs.
+As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
 6 Experiments
-This study primarily conducts comparative experiments on mathematical reasoning tasks.
-The reinforcement learning frameworks included in the comparison are: MindSpeed-RL, which adopts a training–inference shared-resource architecture; the synchronous reinforcement learning version implemented in this work under a training–inference decoupled design; and the asynchronous reinforcement learning version proposed in this study.
-The main objective is to experimentally validate the performance between periodic asynchronous and synchronous training, as well as to compare the proposed periodic asynchronous system with other mainstream synchronous reinforcement learning frameworks on the same platform, while also evaluating the scalability of our periodic asynchronous system.
-The primary evaluation metric is the end-to-end training throughput, such as the average number of tokens trained per second per device (TPSPD).
-This is calculated as the total number of tokens processed over several steps divided by the total elapsed time and then divided by the number of devices used.
-Since no modifications were made to the reinforcement learning algorithm, accuracy metrics are reported only for reference, primarily to verify the correctness of algorithm reproduction.
-The experiments in this study were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight NPUs, each providing 64 GB of memory.
-The intra-node interconnect bandwidth is 196 GB/s, while the inter-node bandwidth is 100 Gb/s.
+Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
+Baseline 0 0.725 0 0 0
+MindSpeed-RL 100 0.733 706.355 198.942 16 61.641
+Sync (ours) 100 0.746 827.306 143.678 16 99.966
+Async (ours) 100 0.758 917.609 82.861 16 192.259
+Table 1: Accuracy and training efficiency comparison on Qwen3-8B.
+All frameworks were trained for 100 steps, and training time and token counts were extracted from logs.
+For our synchronous and asynchronous methods under the training–inference decoupled architecture, 8 NPUs were used for one training instance ($\text{TP}=8,\text{PP}=1$) and 8 NPUs for four inference instances ($\text{TP}=2$).
+MindSpeed-RL used all 16 NPUs for two training instances ($\text{TP}=8,\text{PP}=1$) and eight inference instances ($\text{TP}=2$).
+Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
+Baseline 0 0.726 0 0 0
+MindSpeed-RL 20 0.696 107.874 70.647 64 6.627
+Async (ours) 20 0.737 123.999 21.453 48 33.449
+Table 2: Accuracy and training efficiency comparison on DeepSeek-R1-Distill-Qwen-32B.
+All frameworks were trained for 20 steps, and training time and token counts were taken from logs.
+Our asynchronous setup uses 32 NPUs for one training instance ($\text{TP}=8,\text{PP}=4$) and 16 NPUs for four inference instances ($\text{TP}=4$).
+MindSpeed-RL uses all 64 NPUs for two training instances ($\text{TP}=8,\text{PP}=4$) and 16 inference instances ($\text{TP}=4$).
+This study conducts comparative experiments on mathematical reasoning tasks, evaluating three reinforcement learning frameworks: MindSpeed-RL, which uses a shared-resource architecture for training and inference; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework.
+The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system.
+The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD), calculated as the total number of tokens processed divided by the total elapsed time and the number of devices used.
+Accuracy metrics are provided for reference to verify algorithm correctness, as no modifications were made to the reinforcement learning algorithm.
+Experiments were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight 64 GB NPUs. The intra-node interconnect bandwidth is 196 GB/s, and the inter-node bandwidth is 100 Gb/s.
 6.1 Model and Environment Configuration
-The models compared in this section are Qwen3-8B  [qwen3technicalreport] and DeepSeek-R1-Distill-Qwen-32B  [deepseekai2025deepseekr1incentivizingreasoningcapability] (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.).
-The training data comes from the publicly available math problem dataset DeepScaler [deepscaler2025].
+The models compared in this section are Qwen3-8B  Team (2025) and DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI (2025) (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.).
+The training data comes from the publicly available math problem dataset DeepScaler Luo and others (2025).
 All reinforcement learning experiments use the GRPO algorithm.
-The accuracy test set is the AIME24 [aime24].
+The accuracy test set is the AIME24 Zhang and Math-AI (2024).
 6.2 Comparison between Asynchronous and Synchronous Training
+Given that the prompts for the samples are relatively short while the responses are longer, Shared-Prompt Attention is disabled in the experiments.
 In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k.
-A total of 16 NPUs were used.
-For the synchronous and asynchronous training configurations implemented in this work under the training–inference decoupled architecture, 8 NPUs were allocated to launch a single training instance with tensor parallelism ($\text{TP}=8$) and pipeline parallelism ($\text{PP}=1$).
-The remaining 8 NPUs were used to launch four inference instances with $\text{TP}=2$.
-For MindSpeed-RL, the training configuration used 16 NPUs to launch two training instances ($\text{TP}=8$, $\text{PP}=1$), while all 16 NPUs were also used to launch eight inference instances ($\text{TP}=2$).
-Table 1: Accuracy and training efficiency comparison results on Qwen3-8B.
-Each framework was trained for 100 steps, and the total time consumed and the total number of tokens trained were extracted from the logs.
-Framework GRPO Steps AIME24 Trained Tokens Training Hours NPUs Tokens/Sec/Device
-Baseline 0 0.725 0 0 0 -
-MindSpeed-RL 100 0.733 706.355M 198.942 16 61.641
-Sync (ours) 100 0.746 827.306M 143.678 16 99.966
-Async (ours) 100 0.758 917.609M 82.861 16 192.259
-In terms of TPSPD, the training speed ratio among the three implementations was 1:1.62:3.12.
-For accuracy evaluation, we adopted a rule-based method, where the final answer predicted by the policy model is considered correct if it can be correctly extracted and matches the ground-truth answer in the dataset; otherwise, it is considered incorrect.
-The tests were conducted using the inference engine vLLM 0.9.1.rc, with an output length limit of 32k.
+As shown in Table 1, in terms of TPSPD, the training speed ratio among the three implementations was $\mathbf{1\!:\!1.62\!:\!3.12}$.
+The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments were conducted.
+As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments.
+According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is $\mathbf{1\!:\!5.05}$.
+For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect.
+The tests were conducted using the inference en gine vLLM 0.9.1.rc with an output length limit of 32k.
 Each test sample was evaluated 8 times, with inference parameters set to temperature = 0.7, top_p = 0.95, and top_k = 20.
-In total, 240 test problems were evaluated.
-It is evident that the models trained by the three frameworks achieved similar accuracy on the AIME24 benchmark, with variations within normal random fluctuations, supporting the equivalence in effectiveness between the asynchronous and synchronous methods implemented in this work.
-Figure 5: Training step-wise average reward score comparison
-We also monitored the reward scores throughout the training process of the three frameworks, as depicted in Figure 5.
-The reward score trajectories of our synchronous and asynchronous methods are nearly indistinguishable, further reinforcing the assertion that the proposed asynchronous approach is computationally equivalent to its synchronous counterpart.
-Both curves show a slight upward trend, which can likely be attributed to the baseline model having already been pre-trained on similar data.
-Notably, in the initial phases of training, the reward mean curve for Mindspeed-RL diverges significantly from that of our approach.
-This disparity may arise from differences in system architecture.
-Additionally, it is important to note that, in contrast to our environment, Mindspeed-RL utilizes vLLM version 0.9.1 and vllm_ascend version 0.9.0rc2.
-The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments are conducted.
-For our asynchronous framework, the configuration is as follows: 32 NPUs are used to launch a single training instance with $\text{TP}=8$ and $\text{PP}=4$, while the remaining 16 NPUs are used to launch four inference instances with $\text{TP}=4$.
-In contrast, for the Mindspeed-RL configuration, 64 NPUs are used to launch two training instances with $\text{TP}=8$ and $\text{PP}=4$, and simultaneously, the same 64 NPUs are used to launch 16 inference instances with $\text{TP}=4$.
-Table 2: Accuracy and training efficiency comparison results on DeepSeek-R1-Distill-Qwen-32B.
-Each framework was trained for 20 steps, and the total time consumed and the total number of tokens trained were extracted from the logs.
-Framework GRPO Steps AIME24 Trained Tokens Training Hours NPUs Tokens/Sec/Device
-Baseline 0 0.726 0 0 0 -
-MindSpeed-RL 20 0.679 105.678M 69.116 64 6.636
-Async (ours) 20 0.737 123.999M 21.453 48 33.449
-The evaluation methodology is identical to that used in the first set of experiments.
-According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is 1 : 5.05.
+The results show that the models trained by all three frameworks achieved similar accuracy on the AIME24 benchmark, with any differences falling within the expected range of random fluctuations.
+This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
+We also monitored the reward scores during training for all three frameworks, as shown in Figure 5(a).
+The reward trajectories of our synchronous and asynchronous methods are almost identical, suggesting that the asynchronous approach is computationally equivalent to the synchronous one.
+Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data.
+Figure 5(b) presents a comparison of reward scores between Mindspeed-RL and our asynchronous architecture, and it appears that the trends are generally consistent.
+(a) Synchronous vs. Asynchronous.
+(b) Mindspeed-RL vs. our Asynchronous.
+Figure 5: Training step-wise average reward score comparison.
 Combining the results of the two experimental groups, we obtain a clear comparative plot of the training speeds, as shown in Figure 6.
-Figure 6: TPSPD between MindSpeed-RL and Async (ours)
+Figure 6: TPSPD: MindSpeed-RL vs. Async (ours).
 6.3 Asynchronous Scalability Experiments
 We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. The results are shown in Table 3.
-Table 3: Scalability comparison results. TPSPD performance of training Qwen3-8B on 16, 32, and 64 NPUs using our asynchronous framework.
-Model NPUs Training Instance Inference Instance Training Steps Tokens/Sec/Device
-Qwen3-8B 16 1 4 20 188.162
-32 2 8 20 171.824
-64 4 16 20 163.208
+Model NPUs Training Steps Tokens/Sec/Device
+Qwen3-8B 16 20 188.162
+32 20 171.824
+64 20 163.208
+Table 3: Scalability comparison results. TPSPD performance of training Qwen3-8B on 16, 32, and 64 NPUs using our asynchronous framework. The training–to–inference instance ratio is fixed at 1:4.
+Figure 7: TPS of our framework under different data-parallel scales.
 When expressed in TPS (tokens per second), the results are more intuitively illustrated in Figure 7.
@@
 Overall, the proposed asynchronous reinforcement learning framework demonstrates near-linear scaling in training speed as the number of devices increases.
-Figure 7: TPS of our framework under different scales of data parallelism
 7 Conclusion
-This paper presents an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods.
+This paper introduces an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods.
 The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated.
-Prompt-level asynchronous parallelism maximizes the overlap between inference and training.
-During training, a three-mode (policy, old policy, reference) unified and identically distributed architecture is introduced, along with a group-mask attention mechanism that provides clear advantages when prompts are long and responses are short.
-Experiments show that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform.
-The results also support the equivalence in training effectiveness between the asynchronous and synchronous methods.
-Furthermore, the decoupling of training and inference allows both to be scaled independently and flexibly, achieving optimal balance and maximizing the speed-to-device ratio.
-The decoupling of training architecture and algorithm also implies that the proposed framework can be applied to other on-policy reinforcement learning algorithms without any modifications for compatibility with asynchronous training architectures.
+Prompt-level asynchronous parallelism is employed to maximize the overlap between inference and training.
+A unified three-mode (policy, old policy, reference) architecture with identically distributed components is introduced, along with a shared-prompt attention mechanism that offers distinct advantages for long prompts and short responses.
+Experimental results demonstrate that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform.
+These results further support the equivalence in training effectiveness between the asynchronous and synchronous methods.
+Additionally, the decoupling of training and inference enables independent and flexible scaling, optimizing the balance between speed and device utilization.
+The decoupling of the training architecture and algorithm also ensures the proposed framework’s compatibility with other on-policy reinforcement learning algorithms, allowing them to benefit from the advantages of asynchronous training without requiring algorithmic modifications, nor ensuring the stability of modified algorithms.

### v4

@@
-Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning for Large Language Models
+Periodic Asynchrony: An On-Policy Approach for Accelerating LLM Reinforcement Learning
 Abstract
@@
 It is worth emphasizing that we apply a unified tri-model architecture in the training phase, and we also proposed a shared-prompt attention mask to reduce repetitive computation.
-In practice, these works have achieved at least a threefold overall performance improvement in RL training on NPU platforms, indicating its potential for widespread application.
+In practice, our approach consistently delivers significant end-to-end training efficiency improvements on NPU platforms, indicating its potential for widespread application.
 1 Introduction
@@
 The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation.
-Experiments show that on a 16-NPU setup training an 8B model with 16k context, our approach delivers over a threefold improvement in end-to-end RL throughput.
+Experiments show that when training a 7B-scale model on 16 NPUs, our proposed optimizations achieve over a twofold improvement in end-to-end RL training throughput.
 2 Related Work
@@
 Most recently, AReaL Fu et al.
-(2025) advanced the application of asynchronous RL to LLM reasoning tasks by fully decoupling generation from training.
-It introduces a parameter $\eta$ to control data staleness, enabling better handling of delayed samples.
-To better handle delayed samples, it incorporates a staleness-enhanced PPO variant, effectively modifying the on-policy paradigm to tolerate partially off-policy data.
-Additionally, its training start time depends on the completion of the first inference batch.
-In a similar vein, ROLL Flash Lu et al.
-(2025), also recently introduced, shares conceptual similarities with AReaL.
-Like AReaL, it decouples training from inference and controls the proportion of off-policy data through an async ratio.
-Both AReaL and ROLL Flash are off-policy asynchronous frameworks that improve training efficiency but require modifications to standard on-policy RL algorithms to handle data bias.
-This design trades a small portion of data originally intended for strictly on-policy training to avoid inference-service blocking, effectively shifting it toward off-policy learning.
+(2025) advances asynchronous RL for LLM reasoning by fully decoupling generation from training and introducing explicit staleness control through a parameter $\eta$.
+To tolerate delayed samples, it adopts a staleness-aware PPO variant that relaxes the strict on-policy assumption.
+While effective in improving training efficiency, this design modifies standard on-policy objectives and introduces controlled bias, whose generalization across algorithms and data regimes remains theoretically unclear.
+In parallel, several recent works focus on system-level asynchrony to improve the scalability of LLM post-training.
+ROLL Flash Lu et al.
+(2025), Laminar Sheng et al.
+(2025), and LlamaRL Wu et al.
+(2025) propose fully decoupled asynchronous execution frameworks that eliminate global synchronization barriers and maximize throughput under variable rollout latency.
+Trajectory Balance with Asynchrony Bartoldson et al.
+(2025) further explores decoupling exploration and learning to accelerate post-training.
+These systems primarily address efficiency from a systems perspective and typically rely on standard PPO-style updates, without explicitly constraining policy staleness or preserving strict on-policy correctness.
+For fully asynchronous designs that incorporate off-policy updates, this design effectively trades a small portion of data for off-policy training to prevent transient blocking between inference and learning.
 Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
-In contrast, our method takes a different perspective with several innovations:
+In contrast, our method takes a complementary algorithmic perspective with several innovations:
   • Algorithm-decoupled periodic asynchrony enabling purely on-policy asynchronous acceleration without sacrificing stability.
@@
 15:     $s_{i}\leftarrow$ ProcessSample($p_{i},r_{i}$)
-16:     Compute loss $L(s_{i})$ then backward
+16:     Compute loss $L(s_{i})$ and gradient $\nabla L(s_{i})$
 17:     $O\leftarrow O+\nabla L(s_{i})$
@@
 20:  Move policy’s weights into old-policy’s modules
-21:  Update policy’s weights using gradient $O$
+21:  Update policy’s weights with accumulated gradient $O$
 22:end for
@@
 The decoupled inference–training architecture naturally enables such balancing by allowing independent configuration of inference and training instances.
-Therefore, in theory, when a training batch cannot be processed within a single inference batch, the asynchronous system can achieve more than twice the startup speed of the synchronous system.
+Therefore, in theory, the asynchronous system can achieve nearly twice the end-to-end training speed of the synchronous system.
 5 Implementation
 Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture.
-We incorporate parts of Megatron-Core v0.8.0 and DeepSpeed v0.17.2 Rasley et al.
+We incorporate parts of Megatron-Core and DeepSpeed Rasley et al.
 (2020).
-On the NPU platform, we additionally use MindSpeed v2.1.0_core_r0.8.0, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023).
-For inference, we use vLLM, with ascend_vLLM 0.9.2.rc1 on NPUs.
+On the NPU platform, we additionally use MindSpeed, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023).
+For inference, we use vLLM, with ascend_vLLM on NPUs.
 As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
 6 Experiments
-Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
-Baseline 0 0.725 0 0 0
-MindSpeed-RL 100 0.733 706.355 198.942 16 61.641
-Sync (ours) 100 0.746 827.306 143.678 16 99.966
-Async (ours) 100 0.758 917.609 82.861 16 192.259
-Table 1: Accuracy and training efficiency comparison on Qwen3-8B.
-All frameworks were trained for 100 steps, and training time and token counts were extracted from logs.
-For our synchronous and asynchronous methods under the training–inference decoupled architecture, 8 NPUs were used for one training instance ($\text{TP}=8,\text{PP}=1$) and 8 NPUs for four inference instances ($\text{TP}=2$).
-MindSpeed-RL used all 16 NPUs for two training instances ($\text{TP}=8,\text{PP}=1$) and eight inference instances ($\text{TP}=2$).
-Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
-Baseline 0 0.726 0 0 0
-MindSpeed-RL 20 0.696 107.874 70.647 64 6.627
-Async (ours) 20 0.737 123.999 21.453 48 33.449
-Table 2: Accuracy and training efficiency comparison on DeepSeek-R1-Distill-Qwen-32B.
-All frameworks were trained for 20 steps, and training time and token counts were taken from logs.
-Our asynchronous setup uses 32 NPUs for one training instance ($\text{TP}=8,\text{PP}=4$) and 16 NPUs for four inference instances ($\text{TP}=4$).
-MindSpeed-RL uses all 64 NPUs for two training instances ($\text{TP}=8,\text{PP}=4$) and 16 inference instances ($\text{TP}=4$).
-This study conducts comparative experiments on mathematical reasoning tasks, evaluating three reinforcement learning frameworks: MindSpeed-RL, which uses a shared-resource architecture for training and inference; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework.
+This study conducts comparative experiments on mathematical reasoning tasks, evaluating four reinforcement learning frameworks: MindSpeed-RL, the official NPU training framework, which uses a Megatron backend and adopts a shared-accelerator design for training and inference; VERL, a currently mainstream reinforcement learning framework, which uses an FSDP backend and supports asynchronous execution; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework.
 The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system.
@@
 6.1 Model and Environment Configuration
-The models compared in this section are Qwen3-8B  Team (2025) and DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI (2025) (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.).
-The training data comes from the publicly available math problem dataset DeepScaler Luo and others (2025).
-All reinforcement learning experiments use the GRPO algorithm.
+The models compared in this section are Qwen2.5-7B-Instruct Yang et al.
+(2024), Qwen3-8B Team (2025), and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI (2025), as the current version of MindSpeed-RL does not yet support reinforcement learning training for Qwen3-32B.
+The training data comes from the publicly available math problem datasets DeepScaler Luo and others (2025) and GSM8K Cobbe et al.
+(2021).
+All reinforcement learning experiments use GRPO algorithm.
 The accuracy test set is the AIME24 Zhang and Math-AI (2024).
-6.2 Comparison between Asynchronous and Synchronous Training
-Given that the prompts for the samples are relatively short while the responses are longer, Shared-Prompt Attention is disabled in the experiments.
+6.2 Asynchronous Training Speed Comparison
+Since the prompts in DeepScaler are significantly shorter than the responses, Shared-Prompt Attention is disabled; nevertheless, our method still achieves substantial performance improvements.
 In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k.
-As shown in Table 1, in terms of TPSPD, the training speed ratio among the three implementations was $\mathbf{1\!:\!1.62\!:\!3.12}$.
-The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments were conducted.
+As shown in Table 1, in terms of TPSPD, the training speed ratio among the four implementations was $\mathbf{1\!:\!2.52\!:\!1.62\!:\!3.12}$.
+The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B.
 As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments.
-According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is $\mathbf{1\!:\!5.05}$.
+According to the experimental results, for the 32B-scale model, the TPSPD ratio between MindSpeed-RL and our asynchronous framework reaches $\mathbf{1\!:\!5.05}$ under a 16K context length.
+Due to consistent out-of-memory issues with VERL at 16K, we additionally evaluate VERL under an 8K context length, where the TPSPD ratio between VERL and our asynchronous framework is $\mathbf{1\!:\!1.76}$.
+This experiment also shows that insufficient context-length configuration can lead to a degradation in model accuracy.
+We further observe that, under this regime, inference length is a primary factor affecting end-to-end throughput.
+Given the increased problem difficulty, DeepScaler represents a regime in which inference is more time-consuming than training.
+To study the complementary setting where training dominates inference, we additionally conduct experiments on GSM8K using Qwen2.7-7B-Instruct, where both prompts and responses are relatively short.
+We compare MindSpeed-RL, VERL, and our asynchronous framework under configurations with Shared-Prompt Attention enabled and disabled, and evaluate their end-to-end throughput and accuracy.
+As shown in Table 3, Shared-Prompt Attention significantly reduces the number of training tokens and leads to improved training throughput, with the corresponding TPSPD ratios of $\mathbf{1\!:0.91\!:\!0.27\!:\!2.19}$.
+Setting AIME24 Micro-BS Training Tokens TPSPD
+Base model (8B) 0.725
+MindSpeed-RL 0.733 1 706.355 M 61.641
+VERL (async) 0.746 16 812.084 M 155.521
+Sync (ours) 0.746 1 827.306 M 99.966
+Async (ours) 0.758 1 917.609 M 192.259
+Table 1: All frameworks are trained for 100 steps using 16 NPUs with a global batch size of 32.
+For GRPO, each group performs 32 rollouts.
+Since VERL adopts an FSDP backend, the micro-batch size is set to at least 16 and its asynchronous execution is enabled.
+Setting AIME24 NPUs GBS MBS Training Tokens TPSPD
+Base model (32B) 0.726
+MindSpeed-RL 0.717 64 32 1 107.874 M 6.627
+Async (ours) 0.737 48 32 1 123.999 M 33.449
+VERL (async) 0.617 64 64 64 157.765 M 44.016
+Async (ours) 0.675 64 64 1 188.081 M 77.342
+Table 2: All frameworks are trained for 20 steps.
+For GRPO, each group performs 32 rollouts.
+In the additional group of experiments, the global batch size (GBS) is increased to 64, since VERL requires a minimum micro-batch size (MBS) of 64 when running on 64 NPUs.
+Setting GSM8K Micro-BS Training Tokens TPSPD
+Base model (7B) 0.801
+MindSpeed-RL 0.890 16 72.534 M 199.142
+VERL (async) 0.869 16 90.042 M 182.560
+Async (ours) 0.912 1 84.192 M 54.463
+Async (ours) 0.919 16 60.926 M 435.596
+Table 3: All frameworks are trained on 16 NPUs and complete training on the full GSM8K training set at step 116.
+For GRPO, each group performs 32 rollouts, and the context length is fixed to 1K for all settings.
+In our asynchronous framework, a micro-batch size of 1 indicates that Shared-Prompt Attention is disabled, whereas a micro-batch size of 16 means that 16 rollouts share a single prompt.
+Evaluation is conducted on the GSM8K test set.
 For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect.
-The tests were conducted using the inference en gine vLLM 0.9.1.rc with an output length limit of 32k.
-Each test sample was evaluated 8 times, with inference parameters set to temperature = 0.7, top_p = 0.95, and top_k = 20.
-The results show that the models trained by all three frameworks achieved similar accuracy on the AIME24 benchmark, with any differences falling within the expected range of random fluctuations.
+The tests were conducted using the inference en gine vLLM 0.9.1.rc.
+The results show that the models trained by all four frameworks achieved similar accuracy, with any differences falling within the expected range of random fluctuations.
 This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
-We also monitored the reward scores during training for all three frameworks, as shown in Figure 5(a).
-The reward trajectories of our synchronous and asynchronous methods are almost identical, suggesting that the asynchronous approach is computationally equivalent to the synchronous one.
+We also monitored the reward scores during training for all four frameworks, as shown in Figure 5.
+The reward trajectories of our synchronous and asynchronous methods almost completely overlap, suggesting that the asynchronous approach is computationally equivalent to the synchronous one.
 Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data.
-Figure 5(b) presents a comparison of reward scores between Mindspeed-RL and our asynchronous architecture, and it appears that the trends are generally consistent.
-(a) Synchronous vs. Asynchronous.
-(b) Mindspeed-RL vs. our Asynchronous.
 Figure 5: Training step-wise average reward score comparison.
-Combining the results of the two experimental groups, we obtain a clear comparative plot of the training speeds, as shown in Figure 6.
-Figure 6: TPSPD: MindSpeed-RL vs. Async (ours).
 6.3 Asynchronous Scalability Experiments

### v5

@@
 Abstract
-Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention, with growing efforts to reproduce and apply it.
-However, training efficiency remains a critical challenge.
-In mainstream RL frameworks, inference and training are typically deployed on the same devices.
-While this approach reduces costs through resource consolidation, its synchronous execution imposes a computational coupling that prevents concurrent inference and training.
-In this study, we are returning to the strategy of separating inference and training deployment, and by introducing improvements in the data loader, we transform the conventional synchronous architecture into a periodically asynchronous framework, which allows for demand-driven, independent, and elastic scaling of each component, while the accuracy of the algorithm remains completely equivalent to the synchronization method, with both belonging to the on-policy strategy.
-It is worth emphasizing that we apply a unified tri-model architecture in the training phase, and we also proposed a shared-prompt attention mask to reduce repetitive computation.
-In practice, our approach consistently delivers significant end-to-end training efficiency improvements on NPU platforms, indicating its potential for widespread application.
+Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention for LLM post-training, yet training efficiency remains a critical challenge.
+In mainstream RL frameworks, inference and training are co-located on the same devices, and their synchronous execution prevents concurrent inference and training.
+In this work, we revisit the strategy of separating inference and training deployment, and propose a periodically asynchronous framework that transforms synchronous RL training into an asynchronous producer–consumer pipeline.
+Unlike existing asynchronous approaches that introduce off-policy bias, our design is provably equivalent to its synchronous counterpart, preserving strict on-policy correctness without any algorithmic modifications.
+We further introduce a unified tri-model architecture and a shared-prompt attention mechanism to support efficient asynchronous execution and reduce redundant computation.
+Experiments on NPU platforms demonstrate a three- to five-fold improvement in end-to-end training throughput over mainstream RL frameworks, while maintaining fully comparable accuracy, indicating its potential for widespread application.
 1 Introduction
@@
 The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation.
-Experiments show that when training a 7B-scale model on 16 NPUs, our proposed optimizations achieve over a twofold improvement in end-to-end RL training throughput.
+Experiments on NPU platforms show that our proposed framework achieves a three- to five-fold improvement in end-to-end training throughput over mainstream RL frameworks, with near-linear scaling as the number of devices increases.
 2 Related Work
@@
 These systems primarily address efficiency from a systems perspective and typically rely on standard PPO-style updates, without explicitly constraining policy staleness or preserving strict on-policy correctness.
-For fully asynchronous designs that incorporate off-policy updates, this design effectively trades a small portion of data for off-policy training to prevent transient blocking between inference and learning.
-Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
-In contrast, our method takes a complementary algorithmic perspective with several innovations:
-  • Algorithm-decoupled periodic asynchrony enabling purely on-policy asynchronous acceleration without sacrificing stability.
-  • Decoupling prompt-level inference production from sample-level training consumption, maximizing overlap between the two.
-  • A shared-prompt mask for efficient reuse of forward/backward computation on prompt tokens.
-  • A distribution-consistent design across policy, old-policy, and reference models, simplifying scaling and system structure.
-  • Independent scalability of inference and training instances, preventing bottlenecks in either component.
-Together, these designs enable efficient long-context training with substantially fewer computational resources.
-All components are implemented in project release v5.0.5 (September 5, 2025): https://github.com/janelu9/EasyLLM.
+These fully asynchronous designs effectively trade a degree of on-policy correctness for throughput, preventing transient blocking between inference and learning.
+Although their experiments demonstrate the effectiveness of this approach, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
+In contrast, our method takes a complementary algorithmic perspective: we propose a periodically asynchronous framework that preserves strict on-policy correctness without algorithmic modifications, while introducing a unified tri-model architecture and shared-prompt attention mechanism to further improve system efficiency.
+Unlike existing asynchronous approaches, our design is provably equivalent to synchronous training, offering theoretical guarantees that generalize across on-policy RL algorithms.
 3 Background
 Training RL with Micro-Batching
-It is well known that when using distributed architectures for large-scale model pre-training or fine-tuning, one typically does not train on an entire batch and immediately update the model parameters with the resulting gradients, as is common in single-GPU training for small models.
-A full batch usually contains hundreds or even thousands of samples, and directly processing it in large-model training can easily lead to GPU out-of-memory (OOM) errors.
-Instead, the batch is divided into multiple micro-batches, each containing only one or two samples.
-These micro-batches are processed sequentially, with a buffer tensor used to accumulate the gradients from each micro-step.
-The accumulated gradients are applied to update the model parameters only after all micro-batches within the batch have been processed.
-In practice, these two training strategies are mathematically equivalent.
-The gradient update rule for full-batch training is:
-$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n},$ (1)
-and for micro-batch training, after accumulating gradients over all micro-batches, the update rule is:
-$W_{t+1}=W_{t}-\alpha\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}g_{i,j},$ (2)
-where $M$ is the total number of micro-batches, $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch.
-This demonstrates that both training strategies are mathematically equivalent, but micro-batching offers significant memory savings.
-Consider the objective function of GRPO for generating responses from a single prompt:
-$J_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big).$ (3)
-The PPO-style clipped advantage term can be split into two lines as follows:
+We first establish that micro-batch training is mathematically compatible with on-policy RL objectives, as this forms the theoretical foundation for the gradient accumulation scheme used in our asynchronous framework.
+In distributed large-scale training, micro-batching is a standard technique that partitions a full batch into smaller micro-batches, accumulating gradients across micro-steps before each parameter update.
+This is mathematically equivalent to full-batch training while avoiding out-of-memory errors.
+We extend this equivalence to RL and show it holds for GRPO.
+The GRPO objective for a single prompt with $G$ sampled responses is:
+$J_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big),$ (1)
+where the PPO-style clipped advantage term is defined as:
 $L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})=\min\Bigg($ $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)}A_{i},$
-$\text{clip}\!\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)},1-\varepsilon,\,1+\varepsilon\Big)A_{i}\Bigg).$ (4)
-The KL divergence term is given by:
+$\text{clip}\!\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)},1-\varepsilon,\,1+\varepsilon\Big)A_{i}\Bigg),$ (2)
+and the KL divergence penalty is:
 $D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})=\frac{\pi_{\mathrm{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\frac{\pi_{\mathrm{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-1.$ (3)
-For a batch of $N$ prompts, the objective becomes:
-$J_{\mathrm{batch}}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{\mathrm{GRPO}}^{i}$
-$=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{G}\sum_{j=1}^{G}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big)$
-$=\frac{1}{NG}\sum_{i=1}^{NG}\Big(L_{i}-\beta D_{KL}^{i}\Big).$ (6)
-Define
-$M=\frac{NG}{m}.$ (7)
-Then the batch objective can be rewritten using micro-batches:
-$J_{\mathrm{batch}}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (8)
-This demonstrates that reinforcement learning can also be trained using micro-batches.
-Compared with pre-training or fine-tuning, a batch now contains $NG$ samples, but GPU memory usage is greatly reduced.
-Consequently, reinforcement learning can handle sequences as long as those supported during fine-tuning.
+For a training batch of $N$ prompts, the batch-level objective aggregates over all $NG$ response samples:
+$J_{\mathrm{batch}}=\frac{1}{N}\sum_{i=1}^{N}J_{\mathrm{GRPO}}^{i}=\frac{1}{NG}\sum_{k=1}^{NG}\Big(L_{k}-\beta D_{KL}^{k}\Big).$ (4)
+Partitioning the $NG$ samples into $M=\lfloor NG/m\rfloor$ micro-batches of size $m$, the objective decomposes as:
+$J_{\mathrm{batch}}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big),$ (5)
+which is mathematically equivalent to the full-batch update.
+This confirms that GRPO training is fully compatible with micro-batching: although each batch now contains $NG$ samples rather than $N$, memory consumption remains bounded by the micro-batch size $m$, enabling reinforcement learning over sequences as long as those supported in supervised fine-tuning.
+This property is directly exploited in our asynchronous framework, where samples arriving from the inference queue are accumulated as micro-batches and trained sequentially without waiting for the full batch to complete.
 4 System Design
-4.1 Unified Tri-Model Architecture
+4.1 System Overview
 When performing reinforcement learning with large-scale models using different frameworks for inference and training—for instance, vLLM for inference and Megatron for training—the process generally involves three main steps:
@@
 These models share the same architecture, differing only in their weights.
-In this work, the training phase is implemented using a Megatron-like architecture, where a complete policy model is distributed across multiple computational units according to tensor parallelism and pipeline parallelism scales.
-During the forward computation of the old policy and reference models, each sub-network managed by the policy model is replicated twice across all computational units—once for computing the old policy logits and once for computing the reference logits.
-The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model.
-Figure 1: Triple Models with Shared Distribution.
-The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure 1.
-The three types of models share an identical architecture and are distributed in the same parallel layout.
-This design allows their logits to be computed simultaneously without requiring any additional topology for the reference model.
-The inference stage, in contrast, is handled by an independent inference engine such as vLLM.
-Before training each batch, the weights of the policy model are synchronized to all inference engines.
-The inference engines then receive prompts from the dataset, generate responses, and return them for scoring.
-Each scored response is concatenated with its corresponding prompt to form a training sample, which is fed into all three models for forward computation.
-In a single micro-step, we obtain the logits from the policy, old policy, and reference models simultaneously.
-After all samples in a batch complete their forward passes—and before the policy model’s weights are updated—its current weights are moved into the old policy model’s state_dict.
-4.2 Shared-Prompt Attention
-In reinforcement learning training, each group of samples shares the same prompt.
-In fact, during micro-batch training, if the micro-batch size is greater than 1, the prompt can be reused.
-When the prompt contains many tokens while the generated responses are relatively short, reusing the prompt can significantly reduce redundant computation and GPU memory usage, thereby improving overall training efficiency.
-Compared with standard training, when adopting the shared-prompt approach, the following modifications are made:
-  1. 1.
-    Concatenate the shared prompt with the token IDs of multiple responses. The corresponding labels are concatenated accordingly, while the prompt portion is excluded from the labels.
-    $\left\{\begin{array}[]{l}x=\left[x_{p},x_{r_{1}},x_{r_{2}}\right]\\ y=\left[y_{r_{1}},y_{r_{2}}\right]\end{array}\right$. (9)
-    In this formulation, $x_{p}$ denotes the token IDs of the shared prompt, and $x_{r_{1}},x_{r_{2}}$ correspond to the tokens of the individual responses.
-  2. 2.
-    Rearrange the position indices such that each response starts immediately after the prompt, ensuring positional continuity with the shared prompt.
-    $p=($ $0,1,\dots,|x_{p}|-1,\ |x_{p}|,\dots,|x_{p}|+|x_{r_{1}}|-1,$
-    $|x_{p}|,\dots,|x_{p}|+|x_{r_{2}}|-1)$ (10)
-    Here, the notation $|\cdot|$ indicates sequence length.
-  3. 3.
-    In the attention module, a shared-prompt mask (Figure 2) is used instead of the standard causal mask.
-This design restricts each response token to attend only to the shared prompt, preventing information leakage across responses.
-    Figure 2: Shared-prompt attention mask.
-  4. 4.
-    For probability computation, the prompt tokens are discarded and the loss is computed only over the response tokens:
-    $\pi=-\mathop{\rm CrossEntropy}\nolimits\left(\hat{y}_{|x_{p}|:|x|},y\right),$ (11)
-    where $\hat{y}_{|x_{p}|:|x|}$ represents the predicted logits of the response tokens.
-All other aspects remain the same as in standard training.
-This grouped micro-batch approach is generally more computationally efficient than simply concatenating micro-batches along the first axis, because the prompt tokens only need to be computed once, and padding is not required for shorter samples within a micro-batch.
-The advantage becomes especially pronounced when the prompt is long and the responses are relatively short.
+How these three models are organized and distributed in our system will be detailed in Section 4.2, as their design is closely tied to the requirements of asynchronous execution.
 4.2 Periodic Asynchronous Reinforcement Learning
@@
 This design transforms on-policy reinforcement learning from synchronous to asynchronous execution, significantly improving training efficiency without requiring any modifications to the underlying reinforcement learning algorithm.
+4.2.1 Asynchronous Execution Mechanism
 The reinforcement learning workflow incorporating the temporary sample generator is illustrated in Figure 1. It follows a typical producer–consumer pattern and consists of four main components:
@@
 This queue serves as a buffer that temporarily stores the generated responses from inference service.
-The main process, acting as the consumer, continuously retrieves completed samples from the queue and feeds them to the three models for training.
+The main process, acting as the consumer, continuously retrieves completed samples from the queue and feeds them to the training engine.
 Under this design, the start time of training is determined by the earliest completed sample among all prompts in the batch.
@@
 Therefore, we refer to this design as periodic asynchrony.
+As implied by the consumer-side design, each training step requires the simultaneous computation of three types of logits—policy, old policy, and reference—for every consumed sample.
+To address this, we adopt a unified tri-model architecture in which all three models share an identical parallel layout across computational units, implemented on a Megatron-like backbone distributed according to tensor parallelism and pipeline parallelism scales.
+Each sub-network is replicated twice—once for the old policy logits and once for the reference logits—where the reference network retains the original model weights and the old policy network holds weights delayed by one training step.
+Figure 2: Triple Models with Shared Distribution.
+As illustrated in Figure 2, this shared distribution allows the logits of all three models to be computed simultaneously in a single micro-step.
+After all samples in a batch complete their forward passes, the policy model’s current weights are moved into the old policy model’s state_dict before the next parameter update.
+This design ensures consistent and well-defined weight synchronization at every update boundary, serving as a structural prerequisite for the correctness of the periodic asynchronous framework.
+Furthermore, since all three models share the same parallel topology, their computational resources are managed uniformly, eliminating the need for separate resource allocation and scheduling across model types and further simplifying system design.
+The complete procedure is summarized in Algorithm 1.
 Algorithm 1 Periodic Asynchronous Reinforcement Learning Algorithm
@@
 22:end for
-In this way, the system constitutes a fully on-policy asynchronous reinforcement learning framework that remains decoupled from the underlying RL algorithm, meaning that no algorithmic modifications are required to ensure correct operation within this asynchronous setup.
-Algorithm 1 is the pseudocode for the asynchronous reinforcement learning system proposed in this paper.
+4.2.2 Efficiency Analysis
 (a) Synchronous training step.
@@
 It is clear that when inference speed and training speed are well balanced, the sum of their waiting times is minimized, resulting in optimal end-to-end training throughput on a given hardware configuration.
-The decoupled inference–training architecture naturally enables such balancing by allowing independent configuration of inference and training instances.
 Therefore, in theory, the asynchronous system can achieve nearly twice the end-to-end training speed of the synchronous system.
+Although in rare extreme cases the training time of a single step may be dominated entirely by either inference or training alone, such imbalance can be mitigated through complementary optimizations—for instance, the shared-prompt attention mechanism introduced in Section 4.3 effectively reduces training computation, and the decoupled inference–training architecture naturally enables load balancing by allowing independent scaling of inference and training instances.
+4.2.3 Correctness Analysis
+As shown in Figure 3(b), the asynchronous system trains samples in a different order from the synchronous system, determined by inference completion time rather than the original batch order.
+We now formally establish that neither the on-policy correctness nor the gradient equivalence is compromised by this reordering.
+Proposition 1 (Periodic Weight Consistency). All rollout samples within the same batch are generated by the same policy $\pi_{\theta_{t}}$, i.e.:
+$\forall\,i\in\{1,\ldots,B\}:\quad o_{i}\sim\pi_{\theta_{t}}(\cdot\mid p_{i}).$ (8)
+Proof.
+Line 3 synchronizes all inference workers to $\theta_{t}$ before any rollout begins, and no weight update is performed until Line 21, after all $B$ samples have been consumed.
+Therefore, every $o_{i}$ is sampled exclusively from $\pi_{\theta_{t}}$, satisfying the on-policy condition identically to a synchronous system.
+Having established on-policy sample consistency, we next show that the order in which these samples are consumed during training does not affect the gradient update.
+Proposition 2 (Gradient Permutation Invariance). Let $\sigma$ be any permutation of the $NG$ training samples across micro-batches. The accumulated gradient is invariant to $\sigma$:
+$\nabla_{\theta}J_{\mathrm{batch}}=\frac{1}{NG}\sum_{k=1}^{NG}\nabla_{\theta}\Big(L_{k}-\beta D_{KL}^{k}\Big).$ (9)
+Proof. The double summation $\sum_{i=1}^{M}\sum_{j=1}^{m}$ enumerates all $NG$ samples regardless of their ordering. The result follows directly from the commutativity of finite summation.
+This guarantees that the periodic asynchronous system is algorithmically equivalent to its synchronous counterpart, with any on-policy RL algorithm remaining correct under asynchronous execution without modification.
+Combining Proposition 1 and Proposition 2, we arrive at the main theorem of this section.
+Theorem 1 (Equivalence of Asynchronous and Synchronous Updates). At each training iteration $t$, the periodic asynchronous system produces an identical parameter update to its synchronous counterpart:
+$\Delta\theta_{\mathrm{async}}^{t}=\Delta\theta_{\mathrm{sync}}^{t}.$ (10)
+Consequently, the two systems follow exactly the same training trajectory $\{\theta_{t}\}$, confirming that periodic asynchrony is purely on-policy and algorithmically equivalent to synchronous RL training.
+We can thus conclude that the proposed periodic asynchronous framework constitutes a fully on-policy asynchronous reinforcement learning system that is algorithmically decoupled from the underlying RL algorithm, requiring no modifications to ensure correct operation.
+Its convergence and stability are identical to those of any synchronous on-policy method it is built upon.
+4.3 Shared-Prompt Attention
+In GRPO-based reinforcement learning, all samples within a group are generated from the same prompt, making it possible to share the prompt computation across responses within a micro-batch, which significantly reduces redundant computation and memory consumption when prompts are long relative to responses.
+Compared with standard training, the shared-prompt approach introduces four targeted modifications to the input construction, positional encoding, attention mask, and loss computation respectively:
+  1. 1.
+    Concatenate the shared prompt with the token IDs of multiple responses. The corresponding labels are concatenated accordingly, while the prompt portion is excluded from the labels.
+    $\left\{\begin{array}[]{l}x=\left[x_{p},x_{r_{1}},x_{r_{2}}\right]\\ y=\left[y_{r_{1}},y_{r_{2}}\right]\end{array}\right$. (11)
+    where $x_{p}$ denotes the token IDs of the shared prompt, and $x_{r_{1}},x_{r_{2}}$ correspond to the tokens of the individual responses.
+  2. 2.
+    Rearrange the position indices such that each response starts immediately after the prompt, ensuring positional continuity with the shared prompt.
+    $p=($ $0,1,\dots,|x_{p}|-1,\ |x_{p}|,\dots,|x_{p}|+|x_{r_{1}}|-1,$
+    $|x_{p}|,\dots,|x_{p}|+|x_{r_{2}}|-1)$ (12)
+    where $|\cdot|$ denotes sequence length.
+  3. 3.
+    In the attention module, a shared-prompt mask (Figure 4) is used instead of the standard causal mask.
+This design restricts each response token to attend only to the shared prompt, preventing information leakage across responses.
+    Figure 4: Shared-prompt attention mask.
+  4. 4.
+    For probability computation, the prompt tokens are discarded and the loss is computed only over the response tokens:
+    $\pi=-\mathop{\rm CrossEntropy}\nolimits\left(\hat{y}_{|x_{p}|:|x|},y\right),$ (13)
+    where $\hat{y}_{|x_{p}|:|x|}$ represents the predicted logits of the response tokens.
+Correctness Analysis.
+By construction of the shared-prompt mask (Step 3), each response token attends only to the prompt and its own preceding tokens, with cross-response attention fully masked.
+Therefore, the hidden states and gradients of each response are determined solely by $x_{p}$ and $x_{r_{k}}$, giving:
+$\nabla_{\theta}\mathcal{L}_{\text{shared}}=\sum_{k=1}^{K}\nabla_{\theta}\mathcal{L}_{k},$ (14)
+where $\mathcal{L}_{k}$ denotes the loss on the $k$-th response under standard training.
+This confirms that the shared-prompt approach is mathematically equivalent to standard per-sample training, introducing no approximation or bias into the training objective.
+Complexity Analysis.
+Let $L_{p}$, $L_{r}$, and $K$ denote the prompt length, average response length, and number of responses per micro-batch, respectively.
+In standard training, the prompt tokens are redundantly computed $K$ times, giving a total attention complexity of $\mathcal{O}(K(L_{p}+L_{r})^{2})$.
+Under the shared-prompt approach, the attention computation decomposes into three parts according to the mask structure: prompt self-attention $\mathcal{O}(L_{p}^{2})$, response-to-prompt attention $\mathcal{O}(KL_{r}L_{p})$, and response self-attention $\mathcal{O}(KL_{r}^{2})$, giving a total of $\mathcal{O}(L_{p}^{2}+KL_{r}(L_{p}+L_{r}))$.
+The reduction ratio between the two approaches is:
+$\rho=\frac{L_{p}^{2}+KL_{r}(L_{p}+L_{r})}{K(L_{p}+L_{r})^{2}}$ (15)
+When $L_{p}\gg L_{r}$, $\rho$ approaches $\frac{1}{K}$, yielding an approximately $K$-fold reduction in attention computation.
+Additionally, no padding is required for shorter responses within a micro-batch, further improving hardware utilization.
+These gains are therefore most pronounced in long-prompt, short-response settings, which are common in complex reasoning tasks with GRPO.
 5 Implementation
@@
 As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
+The inference-to-training instance ratio is configurable and set to 1:4 in all experiments, chosen to balance inference and training throughput on the Ascend-910B platform.
+All components described above are publicly available in project release v5.0.5 (September 5, 2025) at https://github.com/janelu9/EasyLLM.
 6 Experiments
-This study conducts comparative experiments on mathematical reasoning tasks, evaluating four reinforcement learning frameworks: MindSpeed-RL, the official NPU training framework, which uses a Megatron backend and adopts a shared-accelerator design for training and inference; VERL, a currently mainstream reinforcement learning framework, which uses an FSDP backend and supports asynchronous execution; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework.
-The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system.
-The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD), calculated as the total number of tokens processed divided by the total elapsed time and the number of devices used.
-Accuracy metrics are provided for reference to verify algorithm correctness, as no modifications were made to the reinforcement learning algorithm.
-Experiments were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight 64 GB NPUs. The intra-node interconnect bandwidth is 196 GB/s, and the inter-node bandwidth is 100 Gb/s.
+We evaluate four reinforcement learning frameworks on mathematical reasoning tasks: MindSpeed-RL, an official NPU training framework using a Megatron backend with shared-accelerator design; VERL, a mainstream framework using an FSDP backend with asynchronous execution; our synchronous baseline under a decoupled training–inference design; and our proposed asynchronous framework.
+The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD).
+Accuracy metrics are provided as reference to verify that no degradation is introduced by our algorithmic design.
 6.1 Model and Environment Configuration
-The models compared in this section are Qwen2.5-7B-Instruct Yang et al.
-(2024), Qwen3-8B Team (2025), and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI (2025), as the current version of MindSpeed-RL does not yet support reinforcement learning training for Qwen3-32B.
+The models compared in this section are Qwen2.5-7B-Instruct Team and others (2024), Qwen3-8B Yang et al.
+(2025), and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI (2025).
 The training data comes from the publicly available math problem datasets DeepScaleR Luo and others (2025) and GSM8K Cobbe et al.
 (2021).
-All reinforcement learning experiments use GRPO algorithm.
-The accuracy test set is the AIME24 Zhang and Math-AI (2024).
-6.2 Asynchronous Training Speed Comparison
-Since the prompts in DeepScaler are significantly shorter than the responses, Shared-Prompt Attention is disabled; nevertheless, our method still achieves substantial performance improvements.
-In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k.
-As shown in Table 1, in terms of TPSPD, the training speed ratio among the four implementations was $\mathbf{1\!:\!2.52\!:\!1.62\!:\!3.12}$.
-The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B.
-As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments.
-According to the experimental results, for the 32B-scale model, the TPSPD ratio between MindSpeed-RL and our asynchronous framework reaches $\mathbf{1\!:\!5.05}$ under a 16K context length.
-Due to consistent out-of-memory issues with VERL at 16K, we additionally evaluate VERL under an 8K context length, where the TPSPD ratio between VERL and our asynchronous framework is $\mathbf{1\!:\!1.76}$.
-This experiment also shows that insufficient context-length configuration can lead to a degradation in model accuracy.
-We further observe that, under this regime, inference length is a primary factor affecting end-to-end throughput.
-Given the increased problem difficulty, DeepScaler represents a regime in which inference is more time-consuming than training.
-To study the complementary setting where training dominates inference, we additionally conduct experiments on GSM8K using Qwen2.7-7B-Instruct, where both prompts and responses are relatively short.
-We compare MindSpeed-RL, VERL, and our asynchronous framework under configurations with Shared-Prompt Attention enabled and disabled, and evaluate their end-to-end throughput and accuracy.
-As shown in Table 3, Shared-Prompt Attention significantly reduces the number of training tokens and leads to improved training throughput, with the corresponding TPSPD ratios of $\mathbf{1\!:0.91\!:\!0.27\!:\!2.19}$.
+All reinforcement learning experiments use the GRPO algorithm.
+The accuracy test set is AIME24 Zhang and Math-AI (2024).
+Experiments were conducted on air-cooled Ascend-910B NPUs, with each node equipped with eight 64 GB NPUs. The intra-node interconnect bandwidth is 196 GB/s, and the inter-node bandwidth is 100 Gb/s.
+Figure 5: Training step-wise average reward score comparison.
+6.2 Training Throughput Comparison
+6.2.1 Comparison with Existing Frameworks
+8B Model on DeepScaleR.
+Table 1 reports results for Qwen3-8B trained on DeepScaleR with a 16K context length.
+Since prompts in this dataset are significantly shorter than responses, Shared-Prompt Attention is disabled for all methods in this experiment.
+Our asynchronous framework achieves a TPSPD of 192.259, outperforming MindSpeed-RL ($\mathbf{3.12\times}$), VERL ($\mathbf{1.24\times}$), and our synchronous baseline ($\mathbf{1.92\times}$).
+The $1.92\times$ speedup over the synchronous baseline is particularly noteworthy, as it closely approaches the theoretical upper bound of $2\times$ derived in Section 4.2, validating the efficiency analysis of the periodic asynchronous framework in a practical setting.
+32B Model on DeepScaleR.
+Table 2 reports results for DeepSeek-R1-Distill-Qwen-32B.
+In the first group, our asynchronous framework running on only 48 NPUs achieves a TPSPD of 33.449, delivering a $\mathbf{5.05\times}$ speedup over MindSpeed-RL running on 64 NPUs.
+This result is particularly significant: not only does our method achieve substantially higher throughput, it does so with fewer hardware resources, demonstrating both efficiency and resource economy.
+In the second group, all frameworks are evaluated at 8K context to align with VERL, which encounters out-of-memory issues at 16K.
+VERL’s accuracy drops sharply to 0.617—well below the base model level—while our framework maintains 0.675, a degradation we attribute to response truncation at insufficient context length.
+Our asynchronous framework achieves a further $\mathbf{1.76\times}$ speedup over VERL even under these favorable conditions for VERL.
+7B Model on GSM8K.
+Table 3 studies the training-dominated regime using Qwen2.5-7B-Instruct on GSM8K with a 1K context.
+Our asynchronous framework with Shared-Prompt Attention enabled achieves a TPSPD of 435.596, the highest among all settings, corresponding to a $\mathbf{2.19\times}$ speedup over MindSpeed-RL and $\mathbf{2.39\times}$ over VERL.
+Notably, both competing frameworks also achieve competitive accuracy, confirming that the throughput gap is not due to shortcuts in training quality.
 Setting AIME24 Micro-BS Training Tokens TPSPD
@@
 Setting AIME24 NPUs GBS MBS Training Tokens TPSPD
-Base model (32B) 0.726
+Base model (32B) 0.696
 MindSpeed-RL 0.717 64 32 1 107.874 M 6.627
-Async (ours) 0.737 48 32 1 123.999 M 33.449
+Sync (ours) 0.725 48 32 1 123.613 M 26.219
+Async (ours) 0.738 48 32 1 123.999 M 33.449
 VERL (async) 0.617 64 64 64 157.765 M 44.016
+Sync (ours) 0.700 64 64 1 185.796 M 46.519
 Async (ours) 0.675 64 64 1 188.081 M 77.342
@@
 In the additional group of experiments, the global batch size (GBS) is increased to 64, since VERL requires a minimum micro-batch size (MBS) of 64 when running on 64 NPUs.
+VERL is evaluated at 8K context due to out-of-memory issues at 16K.
 Setting GSM8K Micro-BS Training Tokens TPSPD
@@
 VERL (async) 0.869 16 90.042 M 182.560
-Async (ours) 0.912 1 84.192 M 54.463
-Async (ours) 0.919 16 60.926 M 435.596
+Async (ours), w/o SPA 0.912 1 84.192 M 54.463
+Sync (ours), w/ SPA 0.930 16 60.586 M 218.396
+Async (ours), w/ SPA 0.919 16 60.926 M 435.596
 Table 3: All frameworks are trained on 16 NPUs and complete training on the full GSM8K training set at step 116.
 For GRPO, each group performs 32 rollouts, and the context length is fixed to 1K for all settings.
-In our asynchronous framework, a micro-batch size of 1 indicates that Shared-Prompt Attention is disabled, whereas a micro-batch size of 16 means that 16 rollouts share a single prompt.
+SPA denotes Shared-Prompt Attention.
+A micro-batch size of 1 indicates SPA is disabled, whereas 16 means 16 rollouts share a single prompt.
 Evaluation is conducted on the GSM8K test set.
-For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect.
-The tests were conducted using the inference en gine vLLM 0.9.1.rc.
-The results show that the models trained by all four frameworks achieved similar accuracy, with any differences falling within the expected range of random fluctuations.
-This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
-We also monitored the reward scores during training for all four frameworks, as shown in Figure 5.
-The reward trajectories of our synchronous and asynchronous methods almost completely overlap, suggesting that the asynchronous approach is computationally equivalent to the synchronous one.
-Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data.
-Figure 5: Training step-wise average reward score comparison.
-6.3 Asynchronous Scalability Experiments
-We also conducted a set of experiments to evaluate the scalability of our framework, using the same configuration as in the first experiment group. The results are shown in Table 4.
+6.2.2 Architectural Analysis
+To isolate the contribution of our architectural design from the benefit of asynchronous execution, we compare Sync (ours) directly against MindSpeed-RL, the strongest competing synchronous framework.
+Under identical synchronous execution, our framework achieves a $\mathbf{1.62\times}$ speedup on the 8B model (Table 1) and a $\mathbf{3.96\times}$ speedup on the 32B model (Table 2), attributable to the efficiency of the decoupled training–inference design and the unified tri-model architecture.
+Furthermore, in the training-dominated GSM8K setting, Sync (ours), w/ SPA (TPSPD: 218.396) already surpasses VERL (async) (TPSPD: 182.560) despite using synchronous execution, demonstrating that Shared-Prompt Attention alone is sufficient to overcome the throughput gap in training-dominated regimes.
+These results confirm that the architectural contributions of this work provide independent and meaningful efficiency gains beyond asynchronous execution alone.
+6.2.3 Ablation Analysis
+Table 3 enables a clean ablation of the two key components proposed in this work: the periodic asynchronous framework and Shared-Prompt Attention. We isolate their respective contributions as follows.
+Effect of Shared-Prompt Attention.
+Comparing Async (ours), w/o SPA (TPSPD: 54.463) with Async (ours), w/ SPA (TPSPD: 435.596) under otherwise identical settings reveals that enabling Shared-Prompt Attention alone yields an $\mathbf{8\times}$ improvement in throughput.
+This gain stems from two sources: a reduction in training tokens from 84.192 M to 60.926 M due to shared prompt computation, and the elimination of padding overhead within micro-batches.
+These results are consistent with the $K$-fold complexity reduction predicted by the analysis in Section 4.3, where $K=16$ rollouts share a single prompt.
+Effect of Periodic Asynchrony.
+Comparing Sync (ours), w/ SPA (TPSPD: 218.396) with Async (ours), w/ SPA (TPSPD: 435.596) isolates the contribution of asynchronous execution under identical architecture and data conditions.
+The asynchronous framework delivers a $\mathbf{2\times}$ speedup, closely matching the theoretical upper bound established in Section 4.2, and confirming that the overlap between inference and training is effectively exploited in practice.
+Combined Effect.
+The two components are complementary in nature: Shared-Prompt Attention reduces the per-step training cost, while periodic asynchrony minimizes idle waiting time between inference and training.
+Together, they deliver a $\mathbf{2.19\times}$ speedup over the strongest competing framework MindSpeed-RL and an $\mathbf{8\times}$ speedup over our unoptimized baseline (Async w/o SPA), demonstrating that the benefits of the two components are largely multiplicative.
+6.2.4 Accuracy Validation
+For accuracy evaluation, we employ a rule-based method where the predicted answer is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise it is deemed incorrect.
+As shown across all tables, the accuracy of our framework remains consistent with competing methods, with differences falling within the expected range of random fluctuation.
+Furthermore, the reward trajectories of our synchronous and asynchronous methods almost completely overlap throughout training, as shown in Figure 5, empirically corroborating the theoretical equivalence established in Theorem 1.
+These results confirm that the substantial throughput gains introduced by our framework come at no cost to training effectiveness.
+6.3 Scalability Analysis
+We conduct a set of experiments to evaluate the scalability of our framework using the same configuration as the first experiment group. The results are shown in Table 4.
 Model NPUs Training Steps Tokens/Sec/Device
@@
 When expressed in TPS (tokens per second), the results are more intuitively illustrated in Figure 6.
-Training with 32 NPUs shows a 1.83× speedup compared to 16 NPUs, while using 64 NPUs results in a 1.9× speedup over 32 NPUs.
-Overall, the proposed asynchronous reinforcement learning framework demonstrates near-linear scaling in training speed as the number of devices increases.
+Training with 32 NPUs shows a $1.83\times$ speedup compared to 16 NPUs, while using 64 NPUs results in a $1.9\times$ speedup over 32 NPUs, demonstrating near-linear scaling in total throughput.
+The moderate decrease in TPSPD with increasing device count is expected, as inter-node communication overhead grows with scale while the training-to-inference ratio is held fixed at 1:4.
+Overall, the proposed asynchronous reinforcement learning framework demonstrates strong scalability as the number of devices increases.
 7 Conclusion
-This paper introduces an asynchronous on-policy reinforcement learning algorithm that maintains the same training stability as synchronous RL methods.
-The feasibility of performing reinforcement learning training with micro-batches is theoretically demonstrated.
-Prompt-level asynchronous parallelism is employed to maximize the overlap between inference and training.
-A unified three-mode (policy, old policy, reference) architecture with identically distributed components is introduced, along with a shared-prompt attention mechanism that offers distinct advantages for long prompts and short responses.
-Experimental results demonstrate that the proposed framework achieves a three- to five-fold improvement in end-to-end performance over current mainstream RL training frameworks on the NPU platform.
-These results further support the equivalence in training effectiveness between the asynchronous and synchronous methods.
-Additionally, the decoupling of training and inference enables independent and flexible scaling, optimizing the balance between speed and device utilization.
-The decoupling of the training architecture and algorithm also ensures the proposed framework’s compatibility with other on-policy reinforcement learning algorithms, allowing them to benefit from the advantages of asynchronous training without requiring algorithmic modifications, nor ensuring the stability of modified algorithms.
+This paper addresses the training efficiency bottleneck in on-policy reinforcement learning by proposing a periodically asynchronous framework that achieves provably equivalent updates to synchronous training.
+By introducing a temporary data generator between the data loader and the trainer, the framework transforms synchronous RL into an asynchronous producer–consumer pipeline, maximizing the overlap between inference and training without any algorithmic modifications.
+We theoretically establish that micro-batch training is fully compatible with on-policy RL objectives, and prove that neither the on-policy correctness nor the gradient equivalence is compromised by asynchronous execution.
+A unified tri-model architecture with identically distributed components is introduced to support efficient and consistent weight synchronization, along with a shared-prompt attention mechanism that significantly reduces redundant computation in long-prompt, short-response settings.
+Experimental results on NPU platforms demonstrate a three- to five-fold improvement in end-to-end throughput over mainstream RL frameworks, while maintaining fully comparable training effectiveness.
+Furthermore, the decoupling of training and inference enables independent and flexible scaling, and the algorithm-agnostic design ensures compatibility with any on-policy RL method, allowing the community to benefit from asynchronous acceleration without sacrificing theoretical guarantees.